### v2

# Decoding Intelligence: A Framework for Certifying Knowledge Comprehension in LLMs

Abstract
Knowledge comprehension capability is an important aspect of human intelligence. As Large Language Models (LLMs) are being envisioned as superhuman agents, it is crucial for them to be proficient at knowledge comprehension. However, existing benchmarking studies do not provide consistent, generalizable, and formal guarantees on the knowledge comprehension capabilities of LLMs. In this work, we propose the first framework to certify knowledge comprehension in LLMs with formal probabilistic guarantees. Our certificates are quantitative — they consist of high-confidence, tight bounds on the probability that a target LLM gives the correct answer on any knowledge comprehension prompt sampled from a distribution. We design and certify novel specifications that precisely represent distributions of knowledge comprehension prompts leveraging knowledge graphs. We certify SOTA LLMs for specifications over the Wikidata5m knowledge graph. We find that the knowledge comprehension capability improves significantly with scaling the size of the models.
1 Introduction
Large Language Models (LLMs) have demonstrated human-level performance for several real-world tasks (Street et al., 2024; Yang et al., 2023b; Bommasani et al., 2022; Harrison, 2024). An important use case for LLMs is knowledge comprehension (Lazaridou et al., 2022; Khattab et al., 2023), that is, they are often used to summarize long texts and webpages (Perplexity, 2023), respond to user queries based on the context (Yang et al., 2018), and serve as adaptive task decomposers for reasoning-based retrieval augmented generation tasks (Yao et al., 2023b). Knowledge comprehension consists of answering questions by extracting relevant information from large, unstructured texts and reasoning with it. Large context windows of millions of tokens in models like Gemini v1.5 (Gemini Team, 2024) reduce reliance on large knowledge corpora of RAG systems and parametric knowledge held by LLMs. Users increasingly provide extensive references during inference to guide responses. This makes analyzing the knowledge-comprehension and reasoning capabilities of popular LLMs crucial. Moreover, knowledge comprehension is considered a basic evaluation of language understanding in human learners, according to the Bloom’s taxonomy (Bloom, 1956). Students are tested on knowledge comprehension tasks at all levels of school education (National Center for Education Statistics, 2024). Standardized tests such as TOEFL (Educational Testing Service, 2024) and IELTS (IDP IELTS, 2024) contain knowledge comprehension as entire assessments. As LLMs are envisioned to become superhuman agents (Xi et al., 2023), it is imperative to formally analyze them on tasks on which humans are extensively tested, like knowledge comprehension.
There are several benchmarking studies on the performance of LLMs for knowledge comprehension (Liang et al., 2023; Chen et al., 2021; Yang et al., 2018; Wang et al., 2023a; Trivedi et al., 2022; Tang & Yang, 2024). Several of these studies use multi-hop question-answering datasets that consist of questions requiring several sequential reasoning steps to obtain the final correct answer. Thus, benchmarking knowledge comprehension often involves analyzing whether the target LLM can combine multiple pieces of information in meaningful ways and reason its way to the correct answer in the prompt, without deviating or hallucinating. However, the empirical nature of prior studies results in inconsistency in their observations (Wei et al., 2023b; Olsson et al., 2022; Shi et al., 2024). Moreover, while they can convey some high-level trends in the performance of popular LLMs, the results are devoid of any formal guarantees on their applicability. Such guarantees are crucial when deploying LLMs in large-scale knowledge-comprehension tasks in critical domains such as medicine or finance, as they give more confidence about the trustworthiness of LLMs before deployment. Our work aims to bridge this gap by introducing a novel formal certification method, QuaCer-C [^1] Quantitative Certification of Knowledge Comprehension, for obtaining certified bounds on the knowledge comprehension capability of LLMs.
Key challenges. We face the following challenges when developing a formal certification framework for knowledge comprehension in LLMs. (1) We need formal representations capturing the knowledge comprehension property, amenable to certification. Such representations should precisely capture large, diverse sets of prompts (created by varying questions, supporting texts, etc.) for knowledge comprehension and their correct responses. (2) Failure examples where the desirable property does not hold are fairly easy to construct for LLMs by appropriate prompt tuning (Xu et al., 2024; Vega et al., 2023), making binary certificates that indicate whether an LLM satisfies a specification trivially false. (3) Giving provable guarantees on LLMs is a hard, open problem, due to the high number of parameters of the models, for which the traditional certifiers (Singh et al., 2019; Shi et al., 2020; Bonaert et al., 2021) would lose significant precision leading to inconclusive analysis. The number of prompts over which we desire the target LLMs to reason correctly is also large, making enumeration-based methods for formal guarantees infeasible.
Our approach. We conceptualize the knowledge comprehension property as a novel formal specification using a knowledge graph. Instead of specifying correctness of LLM responses for all prompts in any given set, we specify a quantitative property, which is the probability of correct response for any knowledge comprehension prompt sampled from a distribution developed using a given knowledge graph. We propose a model-agnostic, black-box quantitative certification approach, QuaCer-C, which circumvents the issues that traditional approaches have with the number of parameters in LLMs and can even work for closed-source API-access LLMs. QuaCer-C generates high-confidence bounds on the quantitative property using queries, leveraging binomial proportion confidence intervals (Clopper & Pearson, 1934). Estimating with bounds is beneficial as they account for the uncertainty in the estimation as well. While formal analysis has been conducted on the individual generations of LLMs in prior work (Quach et al., 2024), there is no analysis for the average-case risk of LLMs in knowledge comprehension. Our certificates contain provable bounds on the probability of getting correct responses for any random knowledge comprehension task sampled from the distributions given in the specifications.
Contributions. We make the following contributions:
  1. 1.
    We specify the knowledge comprehension property desirable from the LLM responses as a formal specification. Our specifications use popular knowledge graphs such as Wikidata5m (Wang et al., 2021) that are augmented with supporting information about each of their entities. The specifications represent a large set of knowledge comprehension prompts with their respective correct answers expected from any target LLM.
  2. 2.
    We model certification in a target LLM as a probability estimation problem and leverage Clopper-Pearson confidence intervals to generate provable, high-confidence bounds on the quantitative property of interest. Our implementation is provided at https://anonymous.4open.science/r/QuaCer_CAnon-4130.
  3. 3.
    We generate the proposed certificates for the popular LLMs: Llama-3, Mistral, Phi-3, GPT-4o, and Gemini-1.5-Pro. We observe that as the number of model parameters increases, the knowledge comprehension capability of the LLM improves. On comparing different model classes, we see Phi-3 models performing the best among the smaller, open-source models.
Our work is the first step towards providing guarantees on the knowledge comprehension and reasoning capabilities of LLMs, to ameliorate the caution needed when using LLMs (Shanahan, 2023) in a systematic way. We anticipate it to go a long way in making LLMs faithful to traditional notions of reasoning, and hence more trustworthy for deployment in critical domains.
2 Background
2.1 Knowledge graph
A knowledge graph ${\mathcal{G}}=({\mathcal{N}},{\mathcal{E}})$ is a collection of nodes ${\mathcal{N}}$ representing entities interconnected by directed edges in ${\mathcal{E}}$ representing their relations (Peng et al., 2023; Ji et al., 2022). They are commonly used in search engines to enhance the relevance of responses to user queries.
Figure 1: A subgraph of Wikidata5m
Hence, major companies develop their own closed-source knowledge graphs. Wikidata5m (Wang et al., 2021), a popular open-source knowledge graph consisting of 5M nodes, is a structured representation of Wikipedia pages. Each Wikidata node corresponds to a Wikipedia page, containing the page’s abstract and a set of aliases that can be synonymously used to refer to the node. Two nodes $(v_{1},v_{2})$, $v_{1},v_{2}\in{\mathcal{N}}$ are connected by a labeled edge if there is a link in the supporting document for $v_{1}$ to that for $v_{2}$. The edge $(v_{1},v_{2})$ is labeled by a set of aliases to describe the relation between the nodes. Figure 1 shows a subgraph of Wikidata5m.
2.2 Large Language Models
Large Language Models (LLMs) are autoregressive causal language models that operate on a vocabulary $\mathcal{V}$, a set of tokens. LLMs takes a sequence of tokens $x_{1},...,x_{n}$ where $x_{i}\in\mathcal{V},n>0$ and outputs a probability distribution over $\mathcal{V}$ for the potential next token $x_{n+1}$. These models are typically pretrained on vast corpora of text data  (Liu et al., 2024) and have shown remarkable capabilities  (Touvron et al., 2023; Gemini Team, 2024; OpenAI, 2024). Numerous benchmarks  (Yang et al., 2018; Rein et al., 2023; Hendrycks et al., 2021) have been developed to evaluate the performance of LLMs on tasks related to multi-step reasoning, knowledge comprehension and question answering. However, there remains a gap in our theoretical understanding of LLMs’ capabilities.
2.3 Information extraction and reasoning
Information extraction (IE) and reasoning are important research problems in natural language processing. IE involves “extracting structured information from unstructured or semi-structured data” (Chen et al., 2022) such as textual documents. Examples of IE are event extraction (Wadden et al., 2019) and relationship extraction (Pawar et al., 2017). Reasoning is the ability of a model to connect multiple facts using the correct logical operations to arrive at a final answer (Huang & Chang, 2023; Zhang et al., 2024). Typically, reasoning capabilities of LLMs are enhanced by using techniques such as Chain of Thought reasoning and its variants (Wei et al., 2023a; Yao et al., 2023a; Wang et al., 2023b), using world models (Hao et al., 2023), etc. It is evaluated in several tasks such as planning (Wang et al., 2024), mathematical reasoning (Imani et al., 2023), commonsense reasoning (Zhao et al., 2023), etc.
3 Certifying knowledge comprehension
Knowledge comprehension is the ability of a model to accurately reason through a multi-hop question (Yang et al., 2018) (combination of multiple simple information extraction questions that should be answered sequentially to arrive at the final answer) and extract the answers to intermediate questions from information provided in its prompt to reach the correct final answer. Thus, knowledge comprehension is a combination of reasoning and information extraction. Figure 2 gives an overview of our certification framework, QuaCer-C. We formally define knowledge comprehension using a knowledge graph ${\mathcal{G}}=({\mathcal{N}},{\mathcal{E}})$ (Section 2.1) next. Our framework is agnostic to the internal structure of the target model ${\mathcal{L}}$ which can be any text-to-text generating model.
  Knowledge Graph Grammar   1. $\gamma$    := $\mathcal{V}^{+}$ 2. $\eta$    := $\mathcal{V}^{+}$ 3. $\mathcal{A}$    := $[\eta_{1},\eta_{2},\dots]$ 4. $v$    := $(\gamma,\mathcal{A})$ 5. $e$    := $((v_{1},v_{2}),\mathcal{A})$ 6. $\mathcal{N}$    := $[v_{1},v_{2},\dots]$ 7. $\mathcal{E}$    := $[e_{1},e_{2},\dots]$ 8. $\mathcal{G}$    := $(\mathcal{N},\mathcal{E})$
Figure 2: Overview of QuaCer-C. (a) A knowledge graph ${\mathcal{G}}$ pivoted on some node $v_{1}$, in this case on ‘Paracetamol Overdose’. (b) A randomly chosen path $\Pi$ originating at $v_{1}$ from the various other possible paths from $v_{1}$ in ${\mathcal{G}}$. (c) A prompt created from $\Pi$ having contexts of the nodes in $\Pi$, a distractor context (highlighted in orange, as the node for ‘Vomiting’ is a distractor for $\Pi$), and a query from $\Pi$. (d) The target LLM’s response to the prompt, validated using the correct answer. (e) Certifier obtains bounds on the probability of correct response using $n$ samples of LLM responses.
3.1 Specifying knowledge comprehension
The adjacent grammar defines a knowledge graph. Let $\mathcal{V}$ denote the vocabulary of ${\mathcal{L}}$’s tokens. $\mathcal{V}^{+}$ denotes the set of concatenation of non-empty sequences of elements of $\mathcal{V}$. Let each node $v$ of ${\mathcal{G}}$ (line 3) consist of a finite list of synonymous names (a.k.a. aliases, $\mathcal{A}$) that can be used to refer to the node, and a context $\gamma$ that provides more information about the node and its relations with other nodes. For example, in Wikidata5m (Wang et al., 2021), each node has aliases consisting of the identifiers mentioned for the subject of the corresponding Wikipedia page and context which is the abstract of the page. Each edge $e$ in ${\mathcal{G}}$ (line 3) is an ordered pair of related nodes where the relation is identified by a set of synonymous aliases $\mathcal{A}$ for the edge. Let $(v_{1},v_{2})$ denote any edge between nodes $v_{1}$ and $v_{2}$ in ${\mathcal{G}}$. We define ${\mathcal{G}}$ (line 3) as a finite collection of nodes ${\mathcal{N}}$ and edges ${\mathcal{E}}$. A path in ${\mathcal{G}}$ (Definition 3.1) is a set of connected nodes in ${\mathcal{N}}$.
Definition 3.1.
(Path in a Knowledge Graph). A path $\Pi=[v_{1},v_{2}\dots,v_{l}]$ is an ordered collection of nodes in a given knowledge graph ${\mathcal{G}}$, where $l>1$, such that $\forall i\in[l-1],v_{i}\in{\mathcal{N}}$, $(v_{i},v_{i+1})\in{\mathcal{E}}$, and $\forall j\in[l],i\neq j\implies v_{i}\neq v_{j}$. $\Pi_{H}\coloneqq v_{1}$ and $\Pi_{T}\coloneqq v_{l}$ are the head and tail nodes respectively of $\Pi$. Let the $i^{th}$ ($i\in[1,l]$) nodes of $\Pi$ from $\Pi_{H}$ and backwards from $\Pi_{T}$ be $\Pi[i]\coloneqq v_{i}$ and $\Pi[-i]\coloneqq v_{l-i+1}$ respectively.
Definition 3.2 describes a multi-hop reasoning problem, derived from a given knowledge graph ${\mathcal{G}}$. As ${\mathcal{G}}$ naturally encodes several multi-hop problems, we use it to form the specification for a target language model ${\mathcal{L}}$, similar to prior works such as (Ho et al., 2020; Jiang et al., 2023b).
Definition 3.2.
(Multi-hop reasoning problem from ${\mathcal{G}}$). Consider any path $\Pi$ (Definition 3.1) of length $l$ in ${\mathcal{G}}$. A multi-hop reasoning problem ${\mathcal{Q}}$ for $\Pi$ is identifying the tail node $\Pi_{T}$, given an alias of its head node $\Pi_{H}$ and aliases of all edges from $\Pi_{H}$ to $\Pi_{T}$ in $\Pi$. Let $v_{\mathcal{A}}$ denote the corresponding aliases of node $v$ in ${\mathcal{G}}$. Let $\mathcal{D}$ be a function that samples a random alias from the given set of aliases.
${\mathcal{Q}}\coloneqq\mathcal{D}(\Pi_{H,\mathcal{A}})\xrightarrow[]{\mathcal{D}((\Pi_{H},\Pi[2])_{\mathcal{A}})}\dots\xrightarrow[]{\mathcal{D}((\Pi[-2],\Pi_{T})_{\mathcal{A}})}?$
${\mathcal{Q}}$ thus involves $l-1$ reasoning steps to get to the final answer, where each step requires correctly identifying intermediate nodes of $\Pi$. Note, however, that correctness of intermediate reasoning steps is generally not evaluated, and accuracy is defined for the final response (Rajpurkar et al., 2016).
To aid ${\mathcal{L}}$ in correctly answering a multi-hop reasoning query ${\mathcal{Q}}$ and reduce hallucination (Dhuliawala et al., 2023), we provide relevant textual information needed to identify the intermediate and final nodes in the prompt. Hence, the overall task of answering ${\mathcal{Q}}$ involves information extraction for identifying intermediate nodes and reasoning to connect the intermediate answers to reach the final answer, which we collectively call knowledge comprehension. Our overall property quantifies the probability of observing correct knowledge comprehension for a random multi-hop reasoning query developed from ${\mathcal{G}}$. We formally define the property for ${\mathcal{L}}$ as a probabilistic program over ${\mathcal{G}}$ in Algorithm 1. We follow the syntax of the imperative probabilistic programming language in (Sankaranarayanan et al., 2013, Figure 3). The language has primitives for sampling from common distributions such as Uniform ($\mathcal{U}$), Bernoulli ($Ber$), etc., and an estimateProbability(.) function that outputs the probability of a random variable attaining a certain value. As all the random sampling steps in Algorithm 1 can operate with any discrete distribution, we use a generic identifier $\mathcal{D}$ (line 1) for samplers of discrete distributions ($\dots$’ denotes samplers for other discrete distributions).
[…]
A prompt for ${\mathcal{L}}$ consists of ${\mathcal{Q}}$ and a context $\Gamma$ containing information relevant to answer ${\mathcal{Q}}$. $\Gamma$ is formed by concatenating ($\odot$) the contexts for all nodes in $\Pi$. Let $v_{\gamma}$ denote the corresponding context of node $v$ in ${\mathcal{G}}$. Prior works (Shi et al., 2023) on analyzing reasoning in LMs have shown the negative influence of irrelevant information (distractor) in prompts on the performance of LMs, which is not ideally expected. Hence, we include distractor texts in $\Gamma$ and specify that the correct response should not be based on the distractor information. The contexts of nodes $\tilde{v}$ adjacent to any node $\Pi[i]$ ($i\in[1,l-2]$) on $\Pi$, such that the relation of $(\Pi[i],\tilde{v})$ is the same as that of $(\Pi[i],\Pi[i+1])$, can serve as effective distractors for ${\mathcal{Q}}$ (Definition 3.3). This is because, at any intermediate step, the model can pick $\tilde{v}$ as the response, which can deviate ${\mathcal{L}}$’s reasoning from $\Pi$. Nodes adjacent to $\Pi[-1]$ and $\Pi[-2]$ are not distractors. For the former, the model must have already reached the final answer before getting to its adjacent nodes, hence, answering ${\mathcal{Q}}$. In the latter, adjacent nodes following same relation are valid correct answers and not distractors. We denote distractor text in $\Gamma$ as the context of randomly sampled nodes from a distribution $\mathcal{D}$ over all distractor nodes of $\Pi$ in ${\mathcal{N}}$. We demonstrate the effects of using distractor text on the performance of SOTA LLMs in Section 4.
Input: ${\mathcal{L}},{\mathcal{G}},v_{1},\rho$
Output: $p$
1:  $\mathcal{D}\coloneqq\mathcal{U}\mid Ber\mid...$
2:  $\Pi\coloneqq[v_{1},v_{2}\coloneqq(\mathcal{D}([v^{\prime}\mid(v_{1},v^{\prime})\in{\mathcal{E}}])),\dots,v_{\mathcal{D}(\{2,\dots,\rho\})}\coloneqq(\mathcal{D}([v^{\prime}\mid(v_{k},v^{\prime})\in{\mathcal{E}}]))]$
3:  ${\mathcal{Q}}\coloneqq\mathcal{D}(\Pi[0]_{\mathcal{A}})\xrightarrow[]{\mathcal{D}(\Pi[0],\Pi[1])_{\mathcal{A}}}\dots\xrightarrow[]{\mathcal{D}(\Pi[-2],\Pi[-1])_{\mathcal{A}}}\textbf{?}$
4:  $\Gamma\coloneqq\texttt{shuffle}([\Pi[0]_{\gamma},\dots,\Pi[-1]_{\gamma},(\mathcal{D}({\mathcal{N}}))_{\gamma},\dots])$
5:  $\mathcal{P}\coloneqq\Gamma\odot{\mathcal{Q}}$
6:  $p\coloneqq\texttt{estimateProbability}({\mathcal{L}}(\mathcal{P})==\texttt{any}(\Pi[-1]_{\mathcal{A}}))$
Algorithm 1 Knowledge comprehension specification
Definition 3.3.
(Distractor node). A distractor node $\tilde{v}$ for a path $\Pi=[v_{1},v_{2},\dots,v_{l}]$ of ${\mathcal{G}}$ is such that $\forall i\in[1,l],\tilde{v}\neq v_{i}$, and $\exists j\in[1,l-2],[(v_{j},\tilde{v})\in{\mathcal{E}}]\wedge[(v_{j},\tilde{v})_{\mathcal{A}}=(v_{j},v_{j+1})_{\mathcal{A}}]$.
Prior works such as (Chen et al., 2024) have shown that LLM performance can vary with information ordering. Hence, we shuffle the information in $\Gamma$ (line 4) to specify that the model’s response should be invariant to the ordering of information. Our final specification in line 6 is the probability that ${\mathcal{L}}$ generates any alias of the last node of the path, which is the correct answer to ${\mathcal{Q}}$. The specification depends on the choices for the different distributions used at various sampling steps, ${\mathcal{G}}$, $v_{1}$, and $\rho$. It leads to certificates for the behavior of ${\mathcal{L}}$ on a given subgraph of ${\mathcal{G}}$ and paths of length at most $\rho$.
3.2 Certification method
Our algorithm certifies the target LLM ${\mathcal{L}}$ by computing an interval $[p_{l},p_{u}]$ containing the value of the probability $p$ (Algorithm 1, line 6) for a given pivot node $v_{1}$ in ${\mathcal{G}}$ with high confidence. We model $p$ as the probability of setting the underlying boolean random variable $\mathcal{R}=({\mathcal{L}}(\mathcal{P})==\texttt{any}(\Pi[-1]_{\mathcal{A}}))$ to true (success). Thus, $\mathcal{R}\sim Ber(p)$. Exactly determining $p$ would require enumerating over all possible $\mathcal{P}$ which can be developed from any path from a subgraph of ${\mathcal{G}}$ with any random aliases, resulting in an infeasible number of possible prompts, as shown in Appendix B.6. Moreover, we want our method to generalize to closed-source LLMs as well, where the internal structures of the models are unknown. Hence, we cannot use any symbolic methods (Mirman et al., 2020) to determine $p$. Thus, to scalably certify the black-box target LM ${\mathcal{L}}$, we estimate $p$ with provably high-confidence (low error) bounds. Confidence is defined as the probability of the true value $p$ being within the bounds, i.e., $Pr[p\in[p_{l},p_{u}]]$. To establish formal guarantees, we want our estimation procedure to be such that the actual confidence is at least the user-specified confidence level, $1-\delta$ (i.e., $Pr[p_{l}\leq p\leq p_{u}]\geq 1-\delta$), where $\delta$ is a small positive constant. Hence we use conservative method of the Clopper-Pearson confidence intervals (Clopper & Pearson, 1934; Brown et al., 2001; Kurz et al., 2014; Collins, 2010), which is known to produce intervals that are guaranteed to have high confidence. To compute the high-confidence bounds on $p$, we make $n$ independent and identically distributed observations of $\mathcal{R}$, out of which we obtain $0\leq k\leq n$ successes. With the $n$ observations and $1-\delta$ confidence, we use Clopper-Pearson confidence intervals to bound $p$.
4 Experiments
4.1 Experimental setup
We certify the following open-source, instruction finetuned (Wei et al., 2022a) models — Llama-3-instruct 8B model (Dubey et al., 2024), Mistral 7B-Instruct-v0.2 (Jiang et al., 2023a), Phi-3 3B and 14B parameter models (Abdin et al., 2024). We also certify 4-bit and 8-bit quantized versions of the open-source models to study the effects of quantization on a model’s knowledge comprehension capabilities. Among the closed-source models with API access, we certify Gemini-1.5 (Gemini Team, 2024) Flash-001 and Pro-002 models and GPT-4o-0827 (OpenAI, 2024).
We use Wikidata5m (Wang et al., 2021) as our knowledge graph after preprocessing (check Appendix B.1.1 for details). To generate challenging and diverse specifications, we sample 50 pivot nodes from two populations: the top 2000 nodes by out-degree in the global graph, and nodes whose subgraph within radius $\rho$ contains at least 2000 vertices. This strategy ensures specifications rooted around any of the pivot nodes have a large number of paths, making enumerative certification (where all possibilities are tested for satisfaction of the specification) impractical. We set the maximum path length parameter as $\rho=5$, as we empirically observe that longer paths could result in queries that are very unrelated to the head node of the path. As our certificates are over all paths with lengths at most $\rho$ in a given subgraph, we equally prioritize the different possible path lengths in $[1,\rho]$, even though paths with longer lengths can be fewer in number than those with shorter lengths. Hence, we define our sampler from our distribution over paths (Algorithm 1, line 2) which first selects a path length from $[1,\rho]$ uniformly randomly and then selects a path having the chosen path length also uniformly randomly. A query is constructed from the selected path by uniformly selecting any aliases for the nodes in the path (Algorithm 1, line 3). We use uniform distributions so as not to bias the results by prioritizing some elements of the underlying sample spaces. Note, however, the distributions can be modified according to the specific certification usecases. We study the variation in the performance of LLMs with varying path lengths within $[1,\rho]$ in Appendix A.
A prompt is constructed for a selected path and query by prepending a context containing the relevant information to answer the query. The design of the context results in $3$ kinds of specifications, which we describe in Section 4.2. We provide some few-shot examples in the context as well, to explain the task to the models. We provide an ablation study on the effects of varying the number of few-shot examples on the performance of the target LLM in Appendix A. We formulate our prompts as multiple choice questions to facilitate evaluation of the LLM responses. The distribution of prompts is primarily determined by the query generation and answer option selection processes, with occasional context adjustments to accommodate model-specific context window constraints. Queries are derived from paths with uniformly distributed lengths, ensuring a balanced representation of reasoning complexity. Answer options are sampled non-uniformly, prioritizing distractors (in the distractor setting), followed by entities from the query path, and finally random entities related to path nodes. This approach balances query complexity and creates challenging answer options. We detail the complete prompt construction process in Appendix B.3. For evaluation of LLM responses, we use a string matching procedure. The detailed design of our evaluator is presented in Appendix B.7. QuaCer-C generates certificates with confidence $1-\delta=95\%$ and number of samples, $n=250$ samples. We conduct our experiments for open-source LLMs on 2 A100 GPUs.
4.2 Different kinds of specifications
We study the certificates for $3$ different kinds of specifications that arise from variations in the construction of context in the prompts to LLMs (Algorithm 1, line 4) — with context shuffling and distractor context (Shuffle Distractor), with context shuffling and without distractor context (Shuffle), and without context shuffling and without distractor context (Vanilla). These settings enable us to study the effects of these operations on the knowledge provided in LLM prompts. When distractor context is provided, it is only for $1$ distractor node, so as to fit the relevant context for the nodes on the considered path within the context windows of the target LLMs. We hypothesize that distractors to nodes later in the path, closer to the tail node which consists of the final answer, would be more challenging for the LLM due to their proximity and similarity to the answer node. Our distractor node sampler (Appendix B.3, Algorithm 4) thus employs a weighted sampling approach to prioritize distractor nodes closer to the path’s tail. We provide an ablation study on the effects of varying the distribution of distractor nodes in Appendix A.
Table 1: Certification results for different LLMs
Model Specification kind Precision Avg. lower bound Avg. upper bound Avg. accuracy
Gemini-1.5-Pro Vanilla - $0.72\pm 0.06$ $0.83\pm 0.05$ $0.78\pm 0.06$
Shuffle - $0.71\pm 0.06$ $0.82\pm 0.05$ $0.77\pm 0.06$
Shuffle Distractor - $0.64\pm 0.09$ $0.75\pm 0.08$ $0.70\pm 0.09$
GPT-4o Vanilla - $0.70\pm 0.06$ $0.81\pm 0.06$ $0.76\pm 0.06$
Shuffle - $0.69\pm 0.06$ $0.80\pm 0.06$ $0.75\pm 0.06$
Shuffle Distractor - $0.62\pm 0.09$ $0.74\pm 0.08$ $0.68\pm 0.09$
Gemini-1.5-Flash Vanilla - $0.46\pm 0.06$ $0.58\pm 0.06$ $0.52\pm 0.06$
Shuffle - $0.45\pm 0.06$ $0.57\pm 0.06$ $0.51\pm 0.06$
Shuffle Distractor - $0.42\pm 0.10$ $0.55\pm 0.10$ $0.48\pm 0.10$
Llama (8B) Vanilla fp16 $0.31\pm 0.09$ $0.43\pm 0.10$ $0.36\pm 0.10$
8bit $0.30\pm 0.05$ $0.42\pm 0.06$ $0.36\pm 0.06$
4bit $0.27\pm 0.05$ $0.39\pm 0.05$ $0.33\pm 0.05$
Shuffle fp16 $0.31\pm 0.05$ $0.44\pm 0.06$ $0.37\pm 0.06$
8bit $0.31\pm 0.06$ $0.44\pm 0.06$ $0.37\pm 0.06$
4bit $0.28\pm 0.07$ $0.40\pm 0.07$ $0.34\pm 0.07$
Shuffle Distractor fp16 $0.30\pm 0.10$ $0.42\pm 0.11$ $0.36\pm 0.11$
8bit $0.30\pm 0.06$ $0.42\pm 0.06$ $0.35\pm 0.06$
4bit $0.25\pm 0.09$ $0.36\pm 0.09$ $0.30\pm 0.09$
Mistral (7B) Vanilla fp16 $0.33\pm 0.05$ $0.45\pm 0.05$ $0.39\pm 0.05$
8bit $0.32\pm 0.05$ $0.44\pm 0.06$ $0.38\pm 0.05$
4bit $0.31\pm 0.06$ $0.43\pm 0.06$ $0.37\pm 0.06$
Shuffle fp16 $0.34\pm 0.05$ $0.46\pm 0.06$ $0.40\pm 0.05$
8bit $0.34\pm 0.06$ $0.46\pm 0.06$ $0.40\pm 0.06$
4bit $0.32\pm 0.05$ $0.44\pm 0.06$ $0.38\pm 0.06$
Shuffle Distractor fp16 $0.33\pm 0.05$ $0.45\pm 0.05$ $0.39\pm 0.05$
8bit $0.33\pm 0.11$ $0.46\pm 0.12$ $0.39\pm 0.12$
4bit $0.28\pm 0.11$ $0.39\pm 0.12$ $0.33\pm 0.11$
Phi-3 (14B) Vanilla fp16 $0.35\pm 0.05$ $0.47\pm 0.05$ $0.41\pm 0.05$
8bit $0.35\pm 0.05$ $0.47\pm 0.05$ $0.41\pm 0.05$
4bit $0.33\pm 0.04$ $0.45\pm 0.05$ $0.39\pm 0.05$
Shuffle fp16 $0.35\pm 0.04$ $0.48\pm 0.04$ $0.41\pm 0.04$
8bit $0.34\pm 0.06$ $0.47\pm 0.06$ $0.40\pm 0.06$
4bit $0.33\pm 0.04$ $0.46\pm 0.05$ $0.39\pm 0.05$
Shuffle Distractor fp16 $0.33\pm 0.11$ $0.45\pm 0.11$ $0.38\pm 0.11$
8bit $0.31\pm 0.08$ $0.43\pm 0.09$ $0.37\pm 0.08$
4bit $0.30\pm 0.08$ $0.42\pm 0.09$ $0.36\pm 0.09$
Phi-3 (3B) Vanilla fp16 $0.34\pm 0.05$ $0.46\pm 0.06$ $0.40\pm 0.05$
8bit $0.32\pm 0.05$ $0.44\pm 0.06$ $0.38\pm 0.06$
4bit $0.32\pm 0.05$ $0.44\pm 0.06$ $0.38\pm 0.06$
Shuffle fp16 $0.34\pm 0.04$ $0.47\pm 0.05$ $0.40\pm 0.05$
8bit $0.32\pm 0.04$ $0.44\pm 0.05$ $0.38\pm 0.05$
4bit $0.32\pm 0.05$ $0.44\pm 0.05$ $0.38\pm 0.05$
Shuffle Distractor fp16 $0.32\pm 0.10$ $0.45\pm 0.10$ $0.38\pm 0.10$
8bit $0.31\pm 0.09$ $0.43\pm 0.10$ $0.37\pm 0.10$
4bit $0.29\pm 0.10$ $0.41\pm 0.10$ $0.35\pm 0.10$
4.3 Certificates
QuaCer-C generates certificates providing high-confidence, tight lower and upper bounds on the probability of a correct LLM response to a random prompt sampled using the prompt constructor from the given distribution of prompts in our specifications. We report the average value of the lower and upper bounds, over our set of specifications that QuaCer-C certifies for each LLM, in Table 1. We also report the average empirical probability (the ratio of correct responses to the total number of prompts, $n$, for each certificate), averaged over the test set. Next, we summarize the key observations from the certification results in Table 1, some of which follow trends from prior works.
Scaling of knowledge comprehension with model size. We observe that the larger models such as Gemini and GPT have significantly higher bounds than those for the smaller models such as Phi-3, Mistral, Llama (Figure 3).
Figure 3: Variation in certification bounds with models (Vanilla, fp16)
The lower bounds of the larger models are higher than the upper bounds of the smaller models, suggesting a paradigm shift in knowledge comprehension capabilities (especially for Gemini-1.5-Pro and GPT-4o). However, as the larger models are also closed-source, we are unaware whether the enhanced knowledge comprehension capabilities could be due to specialized training or finetuning techniques applied on the models. We see that the smaller models have similar certification bounds. Interestingly, Phi3-3B, which is the smallest model we consider, is performing comparatively to its 14B counterparts and also to larger Mistral and Llama models.
Effects of model quantization. We see that with increase in quantization, the model performance deteriorates on knowledge comprehension, which is contrary to prior findings such as Jin et al. (2024) that suggest that 4-bit quantization can retain the model’s knowledge and reasoning capabilities. The drop in the performance is not large, however.
Effects of different kinds of specifications. Our results for the different kinds of specifications — Vanilla, Shuffle, and Shuffle Distractor, indicate that the Vanilla specifications are generally easier for models, leading to higher certification bounds than those for the other specifications. Shuffle Distractor specifications are challenging specifications for all models resulting in consistently lower certification bounds. However, the differences in the bounds’ values are not high across the settings, which could be because of the challenge of large and unstructured contexts from which the model need to identify relevant information in all cases.
Quality of bounding intervals. Table 1 presents the average bounds from the confidence intervals generated for each certificate. A desirable property for the intervals, alongside their high confidence, is that they should be tight, i.e., their range should be small. Tighter intervals indicate precise analysis with small errors. We observe that the average range of the confidence intervals obtained in our experiments is less than $0.12$.
4.4 Case studies
Figure 4: Qualitative analysis of the samples used for certifying knowledge comprehension for the Vanilla specifications on the Wikidata5m subgraph pivoted at the node for ‘Batman Begins’ movie. The context provided in the prompts is not shown for brevity. Wrong model responses are colored red and the correct ones are colored green. The samples are consistent with the certification results, wherein Phi-3 (3B) has lower certification bounds than those for GPT-4o, which has lower bounds than those for Gemini-Pro.
Figure 5: Categorizing predominant types of LLM responses. The LLMs can either do good reasoning or show failure instances caused by distractors or missed reasoning steps.
In this section, we conduct a qualitative analysis of the certification results in Table 1. Firstly, we show the responses of $3$ models in Figure 4 — Phi-3 (3B), GPT-4o, and Gemini — that were obtained when certifying them for the Vanilla specification defined over a subgraph pivoted at the node for ‘Batman Begins’ movie. The samples are representative of the models’ certification results.
Next, we categorize and identify the predominant kinds of model responses and present their examples. In particular, we frequently see the following two failure modes — distracted and missed relation. In the former case, the model gets deviated from the query by following the distractor context in its prompt, resulting in an incorrect answer. In the latter, the model skips some reasoning steps to arrive at the final, correct answer. In other cases of good reasoning, the model accurately follows the steps of the query and arrives at the correct answer. Figure 5 presents examples of the aforementioned kinds of model responses for the GPT-4o model.
5 Related works
In-context learning. As the context windows of LLMs increase (Gemini Team, 2024; Chen et al., 2023; Dubey et al., 2024), more information can be provided to the models to improve their performance. Moreover, few-shot demonstrations (Brown et al., 2020) and examples from related tasks (Qu et al., 2024) can be provided in the context to inform the models about task-specific details, which can result in better performance. This emergent behavior (Wei et al., 2022b; Lu et al., 2024) wherein LLMs can become proficient at a task with few demonstrations provided in their prompts is in-context learning. We use in-context learning to enhance LLMs’ knowledge to answer the prompts and provide few-shot demonstrations for the reasoning steps expected from the models.
Benchmarking LLM intelligence. Several benchmarks have been proposed to study the reasoning (Zhou et al., 2022; Huang & Chang, 2023; Plaat et al., 2024), arithmetic (Yuan et al., 2023; Song et al., 2024; Yang et al., 2023a), planning (Pallagani et al., 2023; Valmeekam et al., 2023; Huang et al., 2022), and question-answering (Yang et al., 2018; Ho et al., 2020; Welbl et al., 2018) capabilities of LLMs, which are integral components of human intelligence. These benchmarks provide empirical insights and trends into the performance of LMs. However, these insights are generally for static datasets and are not guaranteed to generalize. On the other hand, certification methods provide guarantees on, for example, the scope (defined by specifications) and confidence of its claims, as we illustrate in this work.
6 Conclusion and future work
In this work, we present a novel framework to formally certify LLMs for knowledge comprehension. We develop novel specifications that quantify the probability of correct responses over any random knowledge comprehension prompts from a distribution derived from a knowledge graph. Certificates for knowledge comprehension consist of high-confidence bounds on the probability of correct knowledge comprehension, thus providing a method to compare different LLMs with formal guarantees. Our experiments study the variations in knowledge comprehension along the axes of model size, quantization, and task difficulty and present novel trends.
As our work is the first to certify knowledge comprehension in LLMs to the best of our knowledge, there are several possible directions for future work. Our framework can be integrated with knowledge graph construction methods (Ye et al., 2022), to specify and certify LLMs for comprehension and reasoning over less structured/proprietary inputs. The certification method can also be made more sample-efficient. Moreover, our framework can be adapted to domains with other kinds of graphs, such as social network graphs or medical graphs, to certify LLMs for comprehension in domains of social importance.
Appendix A Ablations
A.1 Few Shot Prompts
We conduct an ablation study to examine the impact of varying the number of few-shot examples on Gemini-Flash’s performance in the vanilla task setting. While our default configuration uses two few-shot examples, we extend this analysis to include up to five examples. Interestingly, we observe no significant variation in performance across these different few-shot configurations. The results are presented below in  2.
Table 2: Certification results for different LLMs with different number of few-shot examples
Model Avg. lower bound Avg. upper bound Avg. accuracy
Gemini-1.5-Flash Vanilla (Default) $0.46\pm 0.06$ $0.58\pm 0.06$ $0.52\pm 0.06$
Gemini-1.5-Flash 3Shot Vanilla $0.46\pm 0.06$ $0.58\pm 0.06$ $0.52\pm 0.06$
Gemini-1.5-Flash 4Shot Vanilla $0.46\pm 0.07$ $0.58\pm 0.07$ $0.52\pm 0.07$
Gemini-1.5-Flash 5Shot Vanilla $0.46\pm 0.07$ $0.58\pm 0.07$ $0.52\pm 0.07$
A.2 Distractor Distributions
To assess the impact of distractor distribution on model performance, we implement three distinct distractor distribution strategies:
  1. 1.
    Tail-weighted: Linearly increasing weights towards the tail end of the path, prioritizing distractors near the answer node. This serves as our default setting.
  2. 2.
    Head-weighted: Linearly increasing weights towards the head of the path, emphasizing distractors near the query’s starting point.
  3. 3.
    Uniform: Equal probability of selecting distractors from any position along the path.
We observe no significant differences in either of the settings. The results are presented in  3 below.
Table 3: Certification results for Gemini-Flash with different distractor distributions
Model Avg. lower bound Avg. upper bound Avg. accuracy
Gemini-1.5-Flash Setting 1 (Default) $0.42\pm 0.10$ $0.55\pm 0.10$ $0.48\pm 0.10$
Gemini-1.5-Flash Setting 2 $0.42\pm 0.11$ $0.55\pm 0.11$ $0.48\pm 0.11$
Gemini-1.5-Flash Setting 3 $0.42\pm 0.11$ $0.55\pm 0.11$ $0.48\pm 0.11$
A.3 Model Performances with varying Path Length
Among our certificates, we have queries of various lengths. Here we study the effects on models behavior on queries with varying length by considering the number of hops they require to reason to answer the query(which is 1 less than the path length). To do so, we refer to the number of hops to answer a query as k where $1\leq$ k $<\rho$.
Varying Setting: In figure  6 we show plots for various specifications for the GPT4o model.
Figure 6: Variations in the bounds against the path lengths across various specifications.
Varying Quantization: In figure  7 we show plots when the quantization is varied with the Llama3-8B model on the shuffle specification and their effects on performance.
Figure 7: Variations in the bounds against the path lengths across various quantizations.
Varying Models: In fig  8 we show plots for the shuffle specification and performance across the models(the open-source models use fp16 precision).
Figure 8: Variations in the bounds against the path lengths across various models in the shuffle setting.
In Figure  6, we observe that the performance across settings converges as k increases and the distractor setting is most impactful on the performance for $k=2$.
In Figure  7, we infer that as k increases the performance of the models’ on the task converges across the different quantizations. We hypothesize this is due to the increasing complexity of the reasoning task.
In Figure 8, we see that larger models (GPT-4o, Gemini-Pro) show less severe drop in performance compared to their smaller models. The figure shows that large models may have learnt to better apply 1-step reasoning for multiple steps when compared to their smaller counterparts.
Appendix B Knowledge Graph and Query Generation
This section details our experimental setup for generating multi-hop reasoning queries using the Wikidata5m knowledge graph. We describe the structure of the knowledge graph, the process of generating random paths, formulating queries, and creating answer options including distractors.
B.1 Knowledge Graph Structure
Our experiments are based on the Wikidata5m knowledge graph (KG). The KG has the following key characteristics:
  • Nodes: Each node represents an entity and is associated with a text paragraph from Wikidata5m.
  • Edges: Edges represent relationships between entities.
  • Text Paragraphs: The text associated with each node may contain information relevant to its connected edges.
  • Node and Edge Aliases: Each node and each edge has a set of aliases associated with them which are just different names for them.
This structure allows us to generate queries that require reasoning across multiple hops in the graph.
B.1.1 Preprocessing the wikidata5m knowledge graph
To ensure the generation of unambiguous queries and support the certification process, we preprocess the wikidata5m dataset.
  1. 1.
    Relation Filtering: We remove relations such as ’instance of’, ’subclass of’, and ’part of’ due to their inherent potential for ambiguity in query formulation.
  2. 2.
    Relevant Information Extraction for edges: To ensure the relevance of relationships in the knowledge graph, we require textual evidence for each edge. When entity A is related to entity B, we identify specific sentences in the descriptive text of either entity that explicitly mention any alias of the other entity. We assume these sentences support the relationship’s existence. These sentences are then linked to the edge, providing context that can be used to answer queries about the relationship. This approach ensures that the knowledge graph contains valid relationships and the specific text that justifies them, enhancing the available context for further analysis. If we find no supporting text for an edge, we drop that edge from the knowledge graph.
  3. 3.
    Unicode to ASCII:For consistency within our experiments, we convert all text containing Unicode characters into their respective ASCII approximations.
B.2 Query Generation
We utilize the Wikidata5m knowledge graph for multi-hop query generation. The query generation process involves the following steps:
B.2.1 Random Path Generation
We begin by selecting a pivot node $v_{0}$ in the knowledge graph ${\mathcal{G}}$. From this pivot, we construct a local subgraph ${\mathcal{G}}(v_{0})$ consisting of all paths $\Pi_{v_{0}}$ originating from $v$. This local subgraph serves as the domain for our path generation process. As arbitrary long paths can lose their semantic meaningfulness, we use a constraint $\rho$ to restrict the length of paths from the pivot node in the subgraph to be maximum $\rho$.
Within ${\mathcal{G}}(v_{0})$, we generate a path $\Pi$ using a randomized depth-first search algorithm. The length of this path, denoted as $k_{choice}$, is sampled randomly from the set ${1,2,...,\rho}$ according to a discrete uniform distribution.
This randomized depth-first search traverses the neighbors of each node in ${\mathcal{G}}(v_{0})$ in a random order, which directly corresponds to the sampling process described in line 10 of Algorithm 1. Specifically, at each step, we sample the next node in the path from a discrete uniform distribution over the current node’s neighbors within the local subgraph, expressed as $\sim(\mathcal{D}([v^{\prime}\mid(v_{i},v^{\prime})\in{\mathcal{E}}\land v^{\prime}\in{\mathcal{G}}(v_{0})]))$, where $v_{i}$ is the current node in the path.
To ensure well-defined queries with unique answers, we introduce an additional constraint on path generation. This constraint requires that each generated path be unique in terms of its sequence of relationships. Specifically, traversing the path from the initial node using the specified relations must lead to a single, unambiguous answer node. This approach prevents queries with multiple valid answers, which would complicate the evaluation of the language model’s performance. It’s important to note that this uniqueness constraint applies only to the specific path being generated. Nodes within the path may still have multiple edges with the same relation type to other entities not on the path. This allowance maintains the natural complexity of the knowledge graph structure, where entities can have multiple relationships of the same type with different entities.
The pseudocode for the path generation algorithm is specified in  2
Algorithm 2 Random Path Generation
1:  Input: Graph $G$, Integer $k$, Vertex $source$
2:  Output: $path$
3:  $path\_len\leftarrow$ RandomInteger(1, $k$)
4:  $path\leftarrow$ None
5:  while $path$ is None do
6:     $path\leftarrow$ DFSPath($G$, $source$, $path\_len$)
7:     if not IsUnique($path$then
8:        $path\leftarrow$ None
9:     end if
10:  end while
11:  return  $path$
B.2.2 Query Formulation
Once a valid path $\Pi$ is generated, we convert it into a query string. This process aligns with line 11 in Algorithm 1. The query is constructed by sampling aliases for each node and relation in the path. For example, a path $\Pi=[A,B,C]$ might be converted to a query ”sampled_alias(A) $\rightarrow$ sampled_alias((A, B)) $\rightarrow$ sampled_alias(B) $\rightarrow$ sampled_alias((B, C)) $\rightarrow$ ?”. Here the tuple of two nodes represents their edge. The aliases are sampled randomly from a discrete uniform distribution over the available aliases for a node or an edge.
B.2.3 Example Query Generation
To illustrate our query generation process, consider the scenario of a path in our subgraph as shown in  9.
Figure 9: Potential Path in a Subgraph where Pivot is Chandler Bing
Our algorithm would construct the following query from the path presented in  9:
[…]
This query requires the LLM to reason through two hops in the knowledge graph:
  1. 1.
    Identify the actor who played Chandler Bing (Matthew Perry)
  2. 2.
    Find the birth date of Matthew Perry (19 August 1969)
This example demonstrates how our query generation process creates questions that require multi-hop reasoning, leveraging the structure and relationships within the knowledge graph.
B.3 Prompt Construction
The final prompt is constructed using a template applied to the query. This process involves several steps, each addressing specific requirements:
  • Query Formulation: Convert the generated path into a query string as described earlier.
  • Context: This is the supporting text we provide the LLM to answer the query correctly. We additionally trim the context to fit within the LLM’s context length limits.
  • Few-shot Examples: Include examples to guide the LLM in understanding the query format and expected answer structure.
  • Answer Options Generation: Create a set of possible answers, including the correct one. The LLM has to choose one of these options as the correct one.
  • Distractors: In the distractor setting, we need to find distractors for the query which need to be included in the prompt.
These inclusions ensure that the prompt is comprehensive, fits within model constraints, and provides sufficient guidance for the LLM to generate accurate responses.
B.3.1 Distractor Selection
Distractors are crucial in assessing an LLM’s comprehension and reasoning abilities. We hypothesize that distractors to nodes later in the path closer to the answer would provide more difficulty for the LLM due to their proximity to the answer node. Our distractor selection process, implemented in Algorithm 4, employs a weighted sampling approach to prioritize distractors associated with entities closer to the path’s end. The algorithm first identifies all potential distractors for each node in the path, then assigns weights inversely proportional to their distance from the final answer node. This weighting scheme favors distractors linked to nodes near the path’s end, but still allows for the selection of distractors related to any node in the path. This refers to the sampling procedure in the line 12 of the probabilistic progam specification  1. By performing weighted sampling from this pool, we ensure a balance between highly relevant distractors and a diverse selection across the entire path.
B.3.2 Answer Options
After formulating the query, we generate a set of answer options. This set includes:
  • The correct answer: The last entity in the generated path.
  • Other entities in the path.
  • Related entities: Entities that share some edge with nodes in the path but are not part of the path.
  • Distractors: A distractor is a node in the knowledge graph ${\mathcal{G}}$ that shares a relation with a node in the path, mirroring the relation that continues the path, but the distractor is not itself part of the path. For a formal definition, refer to Definition 3.3. These are only included in the options in the distractor setting.
The process of generating answer options is detailed in Algorithm 3. In the algorithm, we sample answer options from the set described above so we are basically sampling from the nodes as in the probabilistic program specifiction line 12  1. The answer option algorithm assumes that distractors are input in a list according to the order of preference.
Algorithm 3 Generate Answer Options
1:  Input: $correct\_ans$, $distractors$, $path\_entities$, $random\_entities$, $Graph$, $min\_num\_options$
2:  Output: $options$
3:  $options\leftarrow$ [($correct\_ans$] $\cup$ $distractors$
4:  Add path entities to $options$
5:  Add random entities from $random\_entities$ to $options$
[…]
Algorithm 4 Get Best Distractor
1:  Input: Graph $G$, Path $\Pi$
2:  Output: $best\_distractor$
3:  $D\leftarrow$ [] {List of distractors}
4:  $W\leftarrow$ [] {Weights for distractors}
5:  for $i\leftarrow$ 00 to len($\Pi$) $-2$ do
6:     $v\leftarrow\Pi[i]$
7:     $N\leftarrow$ GetNeighbors($G$, $v$)
8:     $N\_distractors\leftarrow$ FilterDistractors($N$, $v$, $\Pi$)
9:     Extend $D$ with $N\_distractors$
10:     Extend $W$ with $[i+1]*len(N\_distractors)$
11:  end for
12:  if $D$ is not empty then
13:     return  WeightedRandomChoice($D$, $W$)
14:  else
15:     return  None
16:  end if
B.4 Context Trimming
To address the input length limitations of various LLMs, we implement a context trimming procedure. Including all text associated with each node in a reasoning path can result in excessively long contexts. Our procedure aims to preserve the most relevant information from the knowledge graph and supporting texts while respecting each model’s maximum input length. This involves identifying relevant sentences per edge in the graph and then trimming the context for each query based on this information.
B.4.1 Finding Relevant Sentences per Edge
Each node in the Wikidata5m knowledge graph has associated textual support for its relations. We utilize this textual information to provide query-relevant context. We need to determine the relevant information from the textual supports for each edge as this would help us trim the contexts accordingly. For each edge $(u,v)$ in the knowledge graph used in the query or answer options generation, we perform the following steps:
  1. 1.
    Collect Aliases and Text: We gather aliases and the associated text paragraphs for both nodes $u$ and $v$.
  2. 2.
    Split into Sentences: We split the text paragraphs of $u$ and $v$ into individual sentences using NLTK.
  3. 3.
    Identify Relevant Sentences: We identify sentences that explicitly link the two nodes. A sentence from $u$’s text is considered relevant if it contains an alias of $v$, and vice versa.
  4. 4.
    Discard Edges without Relevant Sentences: If no relevant sentences are found for an edge, it is deemed unsupported and is discarded from the graph.
  5. 5.
    Prepend First Sentence: To ensure the entity’s primary name or common alias is included, we prepend the first sentence of each node’s text to its list of relevant sentences.
B.4.2 Trimming to Fit Context Length
When constructing the final prompt for the LLM, we prioritize including the most relevant information within the model’s context length limit. Therefore we need to trim the context according to the LLM’s context limit. We use the following procedure (detailed in Algorithm 5):
  1. 1.
    Create Sentence Lists: We create three lists of sentences:
    • $S_{all}$: Contains all sentences from the text paragraphs of nodes involved in the query and answer options.
    • $S_{query}$: Contains all relevant sentences for the edges that constitute the query path.
    • $S_{options}$: Contains all relevant sentences for the edges used to generate the answer options.
  2. 2.
    Construct the Final Context:
    1. (a)
      We prioritize including all sentences from $S_{query}$ as they are directly related to the query.
    2. (b)
      Next, we add as many sentences from $S_{options}$ as possible, given the remaining context length limit.
    3. (c)
      Finally, we fill the remaining space with sentences from $S_{all}$ that have not been included yet, ensuring no sentence is repeated.
Algorithm 5 Context Construction
1:  Input: $S_{all}$, $S_{query}$, $S_{options}$, $L_{max}$
2:  Output: $C_{trimmed}$
3:  $C\leftarrow S_{query}$
4:  ASSERT TokenizedLength($C$) $\leq L_{max}$
5:  $S_{seen}\leftarrow$ UniqueSet($C$)
6:  for each $s$ in $S_{option}$+$S_{all}$ do
7:     if $s\notin S_{seen}$ and TokenizedLength($C$ + $s$) $\leq L_{max}$ then
8:        $C\leftarrow C$ + $s$
9:        Add $s$ to $S_{seen}$
10:     end if
11:  end for
12:  return $C$ as $C_{trimmed}$
[…]
40
B.5 Few-Shot Examples
To guide the LLM towards the desired response format and demonstrate the reasoning process, we include 2 few-shot examples in the prompt. These examples provide a clear illustration of how to approach the multi-hop reasoning task.
We use the following few-shot examples:
Few Shot Examples Common Context: entity_B is the son of entity_A. entity_E is the sister of entity_A. entity_B leads entity_C. Entity_D is a member of Entity_C. Entity_D is a friend of entity_E. entity_E has mother entity_F who likes the services of entity_C. Question 1: entity_A $\rightarrow$ (father of) $\rightarrow$ (leader of) $\rightarrow$ ? Options: 1. entity_F, 2. entity_C, 3. entity_D, 4. entity_E, 5. entity_B Answer: 2. entity_C Explanation: entity_A $\rightarrow$ (father of) entity_B $\rightarrow$ (leader of) entity_C How to get answer: Find who entity_A is father of to get entity_B, then find what B is the leader of to get entity_C. Question 2: entity_B $\rightarrow$ (chief of) $\rightarrow$ (constitutes) $\rightarrow$ (companion of) $\rightarrow$ ? Options: 1. entity_F, 2. entity_C, 3. entity_D, 4. entity_E, 5. entity_A Answer: 4. entity_E Explanation: entity_B $\rightarrow$ (chief of) entity_C $\rightarrow$ (constitutes) entity_D $\rightarrow$ (companion of) entity_E How to get answer: Find what entity_B is the chief of to get entity_C, find what entity_C constitutes to get entity_D, then find the companion of entity_D to get entity_E.
B.6 Final Prompt
The final prompt presented to the LLM is constructed using a template that incorporates several key elements:
Trimmed Context [B.4]: The relevant context extracted and trimmed.
Query [B.2]: The multi-hop query.
Answer Options [B.3.2]: The generated answer options, including the correct answer and distractors.
Few-Shot Examples [B.5]: A set of examples demonstrating the desired response format.
The prompt template is structured as follows:
Prompt Template {few_shot_examples} Actual Query: Given Context: {context} Answer the question: {query} answer the question by selecting the correct answer from the following options: {options} The format for beginning your response is: correct answer: $<option\_number>.\;<answer>,\;because<succinct\;reason>$ follow this exact format and only choose from the given options
Estimating the number of unique prompts: We estimate a lower bound on the number of unique prompts that can be generated from the Wikidata5m Knowledge Graph (KG) by quantifying the potential unique queries within the graph. Each query can be formulated into multiple prompts through variations in answer presentation, thus making query count a conservative estimate. We analyzed the 50 subgraphs employed in our experiments. For each subgraph, we calculated the number of unique paths(upto the maximum path length hyperparameter, $\rho=4$) and calculated the number of possible queries for each path using the number of aliases for each each entity and relation within a path. This analysis provides an estimate of the unique query generation capacity inherent in subgraphs in our KG.
The mean number of unique queries was $3.04\times 10^{15}$ with a median of $1.24\times 10^{15}$. The minimum and maximum observed values were $1.36\times 10^{12}$ and $1.46\times 10^{16}$, respectively.
Importantly, these figures conservatively estimate the number of unique prompts, as they only consider query variations and not the diversity introduced by different answer options. The actual number of unique prompts is likely significantly larger, making exhaustive enumeration of all possible generated prompts infeasible.
B.7 Response Checker function
We implement a simple response checker function to evaluate the correctness of the model’s answers. The function is defined in Algorithm 6. We write a regular expression to account for trivial formatting errors like extra spaces, brackets, incorrect punctuation, etc.
Algorithm 6 Response Checker
1:  Input: $model\_answer$, $correct\_answer\_num$
2:  Output: $is\_correct$
3:  $model\_answer\leftarrow$ LowerCase($model\_answer$)
4:  $correct\_answer\_num\leftarrow$ LowerCase(ToString($correct\_answer\_num$))
5:  $pattern\leftarrow$ SpecializedRegularExpression(”correct answer: ” + $correct\_answer\_num$)
6:  if RegexMatch($pattern$, $model\_answer$then
7:     $is\_correct\leftarrow 1$
8:  else
9:     $is\_correct\leftarrow 0$
10:  end if
11:  return $is\_correct$

### v3

@@
-Decoding Intelligence: A Framework for Certifying Knowledge Comprehension in LLMs
+Certifying Knowledge Comprehension in LLMs
 Abstract
-Knowledge comprehension capability is an important aspect of human intelligence.
-As Large Language Models (LLMs) are being envisioned as superhuman agents, it is crucial for them to be proficient at knowledge comprehension.
-However, existing benchmarking studies do not provide consistent, generalizable, and formal guarantees on the knowledge comprehension capabilities of LLMs.
-In this work, we propose the first framework to certify knowledge comprehension in LLMs with formal probabilistic guarantees.
-Our certificates are quantitative — they consist of high-confidence, tight bounds on the probability that a target LLM gives the correct answer on any knowledge comprehension prompt sampled from a distribution.
-We design and certify novel specifications that precisely represent distributions of knowledge comprehension prompts leveraging knowledge graphs.
-We certify SOTA LLMs for specifications over the Wikidata5m knowledge graph.
-We find that the knowledge comprehension capability improves significantly with scaling the size of the models.
+Large Language Models (LLMs) are increasingly deployed in safety-critical systems where they provide answers based on in-context information derived from knowledge bases.
+As LLMs are increasingly envisioned as superhuman agents, their proficiency in knowledge comprehension—extracting relevant information and reasoning over it to answer questions, a key facet of human intelligence—becomes crucial.
+However, existing evaluations of LLMs on knowledge comprehension are typically conducted on small test sets, but these datasets represent only a tiny fraction of the vast number of possible queries.
+Simple empirical evaluations on these limited test sets raises concerns about the reliability and generalizability of the results.
+In this work, we introduce the first specification and certification framework for knowledge comprehension in LLMs, providing formal probabilistic guarantees for reliability.
+Instead of a fixed dataset, we design novel specifications that mathematically represent prohibitively large probability distributions of knowledge comprehension prompts with natural noise, using knowledge graphs.
+From these specifications, we generate quantitative certificates that offer high-confidence, tight bounds on the probability that a given LLM correctly answers any question drawn from the specification distribution.
+We apply our framework to certify SOTA LLMs in two domains: precision medicine and general question-answering.
+Our results reveal previously unrecognized vulnerabilities in SOTA LLMs due to natural noise in the prompts.
+Additionally, we establish performance hierarchies with formal guarantees among the SOTA LLMs, particularly in the context of precision medicine question-answering.
 1 Introduction
-Large Language Models (LLMs) have demonstrated human-level performance for several real-world tasks (Street et al., 2024; Yang et al., 2023b; Bommasani et al., 2022; Harrison, 2024).
-An important use case for LLMs is knowledge comprehension (Lazaridou et al., 2022; Khattab et al., 2023), that is, they are often used to summarize long texts and webpages (Perplexity, 2023), respond to user queries based on the context (Yang et al., 2018), and serve as adaptive task decomposers for reasoning-based retrieval augmented generation tasks (Yao et al., 2023b).
-Knowledge comprehension consists of answering questions by extracting relevant information from large, unstructured texts and reasoning with it.
-Large context windows of millions of tokens in models like Gemini v1.5 (Gemini Team, 2024) reduce reliance on large knowledge corpora of RAG systems and parametric knowledge held by LLMs.
-Users increasingly provide extensive references during inference to guide responses.
-This makes analyzing the knowledge-comprehension and reasoning capabilities of popular LLMs crucial.
-Moreover, knowledge comprehension is considered a basic evaluation of language understanding in human learners, according to the Bloom’s taxonomy (Bloom, 1956).
-Students are tested on knowledge comprehension tasks at all levels of school education (National Center for Education Statistics, 2024).
-Standardized tests such as TOEFL (Educational Testing Service, 2024) and IELTS (IDP IELTS, 2024) contain knowledge comprehension as entire assessments.
-As LLMs are envisioned to become superhuman agents (Xi et al., 2023), it is imperative to formally analyze them on tasks on which humans are extensively tested, like knowledge comprehension.
-There are several benchmarking studies on the performance of LLMs for knowledge comprehension (Liang et al., 2023; Chen et al., 2021; Yang et al., 2018; Wang et al., 2023a; Trivedi et al., 2022; Tang & Yang, 2024).
-Several of these studies use multi-hop question-answering datasets that consist of questions requiring several sequential reasoning steps to obtain the final correct answer.
-Thus, benchmarking knowledge comprehension often involves analyzing whether the target LLM can combine multiple pieces of information in meaningful ways and reason its way to the correct answer in the prompt, without deviating or hallucinating.
-However, the empirical nature of prior studies results in inconsistency in their observations (Wei et al., 2023b; Olsson et al., 2022; Shi et al., 2024).
-Moreover, while they can convey some high-level trends in the performance of popular LLMs, the results are devoid of any formal guarantees on their applicability.
-Such guarantees are crucial when deploying LLMs in large-scale knowledge-comprehension tasks in critical domains such as medicine or finance, as they give more confidence about the trustworthiness of LLMs before deployment.
-Our work aims to bridge this gap by introducing a novel formal certification method, QuaCer-C [^1] Quantitative Certification of Knowledge Comprehension, for obtaining certified bounds on the knowledge comprehension capability of LLMs.
-Key challenges.
-We face the following challenges when developing a formal certification framework for knowledge comprehension in LLMs.
-(1) We need formal representations capturing the knowledge comprehension property, amenable to certification.
-Such representations should precisely capture large, diverse sets of prompts (created by varying questions, supporting texts, etc.) for knowledge comprehension and their correct responses.
-(2) Failure examples where the desirable property does not hold are fairly easy to construct for LLMs by appropriate prompt tuning (Xu et al., 2024; Vega et al., 2023), making binary certificates that indicate whether an LLM satisfies a specification trivially false.
-(3) Giving provable guarantees on LLMs is a hard, open problem, due to the high number of parameters of the models, for which the traditional certifiers (Singh et al., 2019; Shi et al., 2020; Bonaert et al., 2021) would lose significant precision leading to inconclusive analysis.
-The number of prompts over which we desire the target LLMs to reason correctly is also large, making enumeration-based methods for formal guarantees infeasible.
+Large Language Models (LLMs) have demonstrated human-level performance in various real-world tasks (Street et al., 2024; Yang et al., 2023; Bommasani et al., 2022; Harrison, 2024).
+A notable application of LLMs is in question-answering systems, including chatbots, which are prompted with relevant, but unstructured information from knowledge bases to generate accurate answers—an ability we refer to as knowledge comprehension.
+Knowledge comprehension is a crucial aspect of language understanding, typically assessed in human learners across all education levels (Bloom, 1956; National Center for Education Statistics, 2024; Educational Testing Service, 2024; IDP IELTS, 2024).
+However, despite being envisioned as superhuman agents (Xi et al., 2023), SOTA LLMs often struggle with basic knowledge comprehension tasks in real-world applications.
+For example, in the adjacent figure, Gemini-1.5-Flash gets confused with the in-context information and provides an incorrect and potentially fatal response in a critical scenario.
+Given the widespread relevance of knowledge comprehension tasks, numerous benchmarking studies have been conducted to evaluate LLMs in this domain (Liang et al., 2023; Chen et al., 2021; Yang et al., 2018; Wang et al., 2023; Tang & Yang, 2024).
+However, their results have been inconsistent.
+For example, Wei et al.
+(2023) demonstrates that larger models are robust to label noise in test sets, while Shi et al.
+(2024) argues that such models can become distracted by this noise.
+Conversely, Olsson et al.
+(2022) indicates that models exhibit similar induction heads for in-context learning, regardless of their size.
+These discrepancies can be attributed to the following inherent limitations in the nature of these evaluations.
+Firstly, LLMs are trained on all available texts, including the standard benchmarking datasets.
+As a result, models may have encountered the exact prompts used in some test sets, leading to test set leakage during evaluation (Mirzadeh et al., 2024).
+Consequently, evaluations relying on these test sets become unreliable.
+Secondly, even in cases where the benchmark studies construct their own datasets, these evaluations remain empirical and assess only a limited subset of potential inputs from the vast space of possible knowledge comprehension prompts.
+Thirdly, for deployment in practical scenarios, it is also crucial for the evaluations to be comprehensive and reliable - potentially with a guarantee for their validity.
+Therefore, the following question arises: How can we assess LLM performance in knowledge comprehension tasks, ensuring reliability of accuracy and generalizability?
+We argue that to reliably deploy LLMs in safety-critical applications, their empirical evaluation on limited test sets is not sufficient.
+Instead, they must be evaluated on diverse and large sets of prompts (consisting of different querying styles, large in-context information that can distract models like above, noise, etc.) for knowledge comprehension.
+Further instead of naively creating prohibitively large test sets, we propose to use precise mathematical representations of these prompts, aka specifications, to enable a computationally tractable evaluation of LLMs.
+For reliability of the evaluations, it is desirable to have guarantees on their validity.
+Although prior works (Seshia et al., 2018; Singh et al., 2019; Shi et al., 2020; Bonaert et al., 2021) have proposed formal specifications and certification with guarantees as a method to evaluate traditional neural networks, we cannot directly use them.
+This is because there exist no formal specifications for knowledge comprehension.
+Also, certifying LLMs is hard due to a high number of model parameters and specialized nonlinearities, for which traditional certifiers would lose significant precision, leading to inconclusive analysis.
 Our approach.
-We conceptualize the knowledge comprehension property as a novel formal specification using a knowledge graph.
-Instead of specifying correctness of LLM responses for all prompts in any given set, we specify a quantitative property, which is the probability of correct response for any knowledge comprehension prompt sampled from a distribution developed using a given knowledge graph.
-We propose a model-agnostic, black-box quantitative certification approach, QuaCer-C, which circumvents the issues that traditional approaches have with the number of parameters in LLMs and can even work for closed-source API-access LLMs.
-QuaCer-C generates high-confidence bounds on the quantitative property using queries, leveraging binomial proportion confidence intervals (Clopper & Pearson, 1934).
-Estimating with bounds is beneficial as they account for the uncertainty in the estimation as well.
-While formal analysis has been conducted on the individual generations of LLMs in prior work (Quach et al., 2024), there is no analysis for the average-case risk of LLMs in knowledge comprehension.
-Our certificates contain provable bounds on the probability of getting correct responses for any random knowledge comprehension task sampled from the distributions given in the specifications.
-Contributions. We make the following contributions:
+Our work proposes a novel formal specification and certification framework, LLMCert-C (LLM Certification of Knowledge Comprehension) that provides reliable performance assessment of the knowledge comprehension capabilities of LLMs.
+To encode meaningful and diverse knowledge comprehension prompts, we rely on a structured representation of ground truth knowledge - a knowledge graph.
+This representation enables us to define new probabilistic specifications, where each specification represents a distribution over knowledge comprehension prompts.
+The prompt distributions, derived from knowledge graphs, are generated with prohibitively large support, consisting of practical prompts with queries over the knowledge graph.
+These prompts contain substantial amounts of unstructured, relevant information, along with naturally occurring noise—such as distracting texts and information shuffling—that can influence the LLM’s responses.
+A certificate for a given specification quantifies the probability that the LLM provides a correct response to any prompt sampled from the specification distribution.
+LLMCert-C employs a quantitative certification approach that overcomes the limitations of traditional methods for LLM architectures by using a query-based method relying on input-output examples.
+This also enables using LLMCert-C for closed-source, API-accessible models.
+In contrast to quantitative certification, binary certificates (Gehr et al., 2018; Wang et al., 2021a) for specification compliance can be trivially invalidated due to the ease of constructing failure examples where the desirable property does not hold (Xu et al., 2024; Vega et al., 2023).
+LLMCert-C generates probabilistic formal guarantees consisting of high-confidence bounds on the probability of correct knowledge comprehension.
+LLMCert-C leverages binomial proportion confidence intervals (Clopper & Pearson, 1934) to generate high-confidence, tight certification bounds to estimate the underlying probability measure.
+The use of bounds is particularly advantageous as they also account for the inherent uncertainty in the estimation.
+We certify with 2 practical knowledge graphs - PrimeKG (Chandak et al., 2023) over precision medicine knowledge and Wikidata5m (Wang et al., 2021b) over general knowledge.
+We use the generated certificates to establish novel performance hierarchies with formal guarantees among SOTA LLMs.
+We also show a consistent decline in LLM performance due to noise in prompts.
+While formal analysis has been conducted on individual generations of LLMs (Quach et al., 2024) and on counterfactual bias (Chaudhary et al., 2024) in prior work, there is no analysis for the average-case performance of LLMs in knowledge comprehension.
+Figure 1 gives an overview of LLMCert-C.
+Main Contributions:
   1. 1.
-    We specify the knowledge comprehension property desirable from the LLM responses as a formal specification.
-Our specifications use popular knowledge graphs such as Wikidata5m (Wang et al., 2021) that are augmented with supporting information about each of their entities.
-The specifications represent a large set of knowledge comprehension prompts with their respective correct answers expected from any target LLM.
+    We define knowledge comprehension desirable from LLM responses as a formal specification, using knowledge graphs.
+We develop novel specifications over large sets of prompts with natural noise, such as distracting text and shuffled information order.
+We generate and experiment with specifications in safety-critical applications in precision medicine and general question-answering using popular knowledge graphs such as PrimeKG and Wikidata5m.
   2. 2.
-    We model certification in a target LLM as a probability estimation problem and leverage Clopper-Pearson confidence intervals to generate provable, high-confidence bounds on the quantitative property of interest.
-Our implementation is provided at https://anonymous.4open.science/r/QuaCer_CAnon-4130.
+    We model certification of any target LLM with query-access for a given specification as a probability estimation problem and leverage confidence intervals to generate high-confidence guaranteed bounds on the probability of correct knowledge comprehension.
+We open-source our code at https://github.com/uiuc-focal-lab/LLMCert-C and provide a note to practitioners in Appendix A.
   3. 3.
-    We generate the proposed certificates for the popular LLMs: Llama-3, Mistral, Phi-3, GPT-4o, and Gemini-1.5-Pro.
-We observe that as the number of model parameters increases, the knowledge comprehension capability of the LLM improves.
-On comparing different model classes, we see Phi-3 models performing the best among the smaller, open-source models.
-Our work is the first step towards providing guarantees on the knowledge comprehension and reasoning capabilities of LLMs, to ameliorate the caution needed when using LLMs (Shanahan, 2023) in a systematic way.
-We anticipate it to go a long way in making LLMs faithful to traditional notions of reasoning, and hence more trustworthy for deployment in critical domains.
-2 Background
-2.1 Knowledge graph
-A knowledge graph ${\mathcal{G}}=({\mathcal{N}},{\mathcal{E}})$ is a collection of nodes ${\mathcal{N}}$ representing entities interconnected by directed edges in ${\mathcal{E}}$ representing their relations (Peng et al., 2023; Ji et al., 2022).
-They are commonly used in search engines to enhance the relevance of responses to user queries.
-Figure 1: A subgraph of Wikidata5m
-Hence, major companies develop their own closed-source knowledge graphs.
-Wikidata5m (Wang et al., 2021), a popular open-source knowledge graph consisting of 5M nodes, is a structured representation of Wikipedia pages.
-Each Wikidata node corresponds to a Wikipedia page, containing the page’s abstract and a set of aliases that can be synonymously used to refer to the node.
-Two nodes $(v_{1},v_{2})$, $v_{1},v_{2}\in{\mathcal{N}}$ are connected by a labeled edge if there is a link in the supporting document for $v_{1}$ to that for $v_{2}$.
-The edge $(v_{1},v_{2})$ is labeled by a set of aliases to describe the relation between the nodes.
-Figure 1 shows a subgraph of Wikidata5m.
-2.2 Large Language Models
-Large Language Models (LLMs) are autoregressive causal language models that operate on a vocabulary $\mathcal{V}$, a set of tokens.
-LLMs takes a sequence of tokens $x_{1},...,x_{n}$ where $x_{i}\in\mathcal{V},n>0$ and outputs a probability distribution over $\mathcal{V}$ for the potential next token $x_{n+1}$.
-These models are typically pretrained on vast corpora of text data  (Liu et al., 2024) and have shown remarkable capabilities  (Touvron et al., 2023; Gemini Team, 2024; OpenAI, 2024).
-Numerous benchmarks  (Yang et al., 2018; Rein et al., 2023; Hendrycks et al., 2021) have been developed to evaluate the performance of LLMs on tasks related to multi-step reasoning, knowledge comprehension and question answering.
-However, there remains a gap in our theoretical understanding of LLMs’ capabilities.
-2.3 Information extraction and reasoning
-Information extraction (IE) and reasoning are important research problems in natural language processing.
-IE involves “extracting structured information from unstructured or semi-structured data” (Chen et al., 2022) such as textual documents.
-Examples of IE are event extraction (Wadden et al., 2019) and relationship extraction (Pawar et al., 2017).
-Reasoning is the ability of a model to connect multiple facts using the correct logical operations to arrive at a final answer (Huang & Chang, 2023; Zhang et al., 2024).
-Typically, reasoning capabilities of LLMs are enhanced by using techniques such as Chain of Thought reasoning and its variants (Wei et al., 2023a; Yao et al., 2023a; Wang et al., 2023b), using world models (Hao et al., 2023), etc.
-It is evaluated in several tasks such as planning (Wang et al., 2024), mathematical reasoning (Imani et al., 2023), commonsense reasoning (Zhao et al., 2023), etc.
+    We observe that as model parameters increase, knowledge comprehension in LLMs improves with high confidence.
+We show a consistent decline in SOTA LLMs’ performance at answering practical prompts with natural noise.
 2 Certifying knowledge comprehension
-Knowledge comprehension is the ability of a model to accurately reason through a multi-hop question (Yang et al., 2018) (combination of multiple simple information extraction questions that should be answered sequentially to arrive at the final answer) and extract the answers to intermediate questions from information provided in its prompt to reach the correct final answer.
-Thus, knowledge comprehension is a combination of reasoning and information extraction.
-Figure 2 gives an overview of our certification framework, QuaCer-C.
-We formally define knowledge comprehension using a knowledge graph ${\mathcal{G}}=({\mathcal{N}},{\mathcal{E}})$ (Section 2.1) next.
-Our framework is agnostic to the internal structure of the target model ${\mathcal{L}}$ which can be any text-to-text generating model.
-  Knowledge Graph Grammar   1.
+Knowledge comprehension is the ability of a model to retrieve and reason over relevant information from context or parametric knowledge to answer questions succinctly and correctly, according to the provided context.
+Figure 1 gives an overview of our knowledge comprehension certification framework, LLMCert-C.
+Our framework is agnostic to the internal structure of the target model ${\mathcal{L}}$ allowing any text-to-text generating model.
+Figure 1: Overview of LLMCert-C.
+(a) A KG ${\mathcal{G}}$.
+(b) A randomly chosen path DAG ${\mathcal{G}}^{\prime}$ originating at fixed node ‘Paracetamol Overdose’, from the various other possible DAGs from node in ${\mathcal{G}}$.
+(c) A prompt created from ${\mathcal{G}}^{\prime}$ having contexts of the nodes in ${\mathcal{G}}^{\prime}$, a distractor context (highlighted in orange, as the node for ‘Vomiting’ is a distractor for ${\mathcal{G}}^{\prime}$), and a query from ${\mathcal{G}}^{\prime}$.
+(d) The target LLM’s response to the prompt, validated using the correct answer.
+(e) Certifier obtains bounds on the probability of correct response using $n$ samples of LLM responses to randomly sampled prompts.
+2.1 Specifying knowledge comprehension
+We formally define knowledge comprehension using a knowledge graph (KG, §B.1) ${\mathcal{G}}=({\mathcal{N}},{\mathcal{E}})$.
+The adjacent grammar defines a KG.
+Let $\mathcal{V}$ denote the vocabulary of ${\mathcal{L}}$’s tokens.
+$\mathcal{V}^{+}$ denotes the set of concatenation of non-empty sequences of elements of $\mathcal{V}$.
+  KG Grammar   1.
 $\gamma$    := $\mathcal{V}^{+}$ 2.
@@
 $\mathcal{G}$    := $(\mathcal{N},\mathcal{E})$
-Figure 2: Overview of QuaCer-C.
-(a) A knowledge graph ${\mathcal{G}}$ pivoted on some node $v_{1}$, in this case on ‘Paracetamol Overdose’.
-(b) A randomly chosen path $\Pi$ originating at $v_{1}$ from the various other possible paths from $v_{1}$ in ${\mathcal{G}}$.
-(c) A prompt created from $\Pi$ having contexts of the nodes in $\Pi$, a distractor context (highlighted in orange, as the node for ‘Vomiting’ is a distractor for $\Pi$), and a query from $\Pi$.
-(d) The target LLM’s response to the prompt, validated using the correct answer.
-(e) Certifier obtains bounds on the probability of correct response using $n$ samples of LLM responses.
-3.1 Specifying knowledge comprehension
-The adjacent grammar defines a knowledge graph.
-Let $\mathcal{V}$ denote the vocabulary of ${\mathcal{L}}$’s tokens.
-$\mathcal{V}^{+}$ denotes the set of concatenation of non-empty sequences of elements of $\mathcal{V}$.
-Let each node $v$ of ${\mathcal{G}}$ (line 3) consist of a finite list of synonymous names (a.k.a.
+Let each node $v$ of ${\mathcal{G}}$ (line 2.1) consist of a finite list of synonymous names (a.k.a.
 aliases, $\mathcal{A}$) that can be used to refer to the node, and a context $\gamma$ that provides more information about the node and its relations with other nodes.
-For example, in Wikidata5m (Wang et al., 2021), each node has aliases consisting of the identifiers mentioned for the subject of the corresponding Wikipedia page and context which is the abstract of the page.
-Each edge $e$ in ${\mathcal{G}}$ (line 3) is an ordered pair of related nodes where the relation is identified by a set of synonymous aliases $\mathcal{A}$ for the edge.
-Let $(v_{1},v_{2})$ denote any edge between nodes $v_{1}$ and $v_{2}$ in ${\mathcal{G}}$.
-We define ${\mathcal{G}}$ (line 3) as a finite collection of nodes ${\mathcal{N}}$ and edges ${\mathcal{E}}$.
-A path in ${\mathcal{G}}$ (Definition 3.1) is a set of connected nodes in ${\mathcal{N}}$.
+Let $v_{\mathcal{A}}$ and $v_{\gamma}$ denote the aliases and context of $v$ respectively.
+For example, in Wikidata5m (Wang et al., 2021b), each node has aliases consisting of the identifiers mentioned for the subject of the corresponding Wikipedia page and context which is the abstract of the page.
+Each (directed) edge $e$ in ${\mathcal{G}}$ (line 2.1) is an ordered pair of related nodes where the relation is identified by a set of synonymous aliases $\mathcal{A}$ for the edge.
+Let $(v_{1},v_{2})$ denote any edge from node $v_{1}$ to $v_{2}$ in ${\mathcal{G}}$, where $(v_{1},v_{2})_{\mathcal{A}}$ denotes the aliases of the edge.
+${\mathcal{G}}$ (line 2.1) is a finite collection of nodes ${\mathcal{N}}$ and edges ${\mathcal{E}}$.
+A directed acyclic sub-graph (DAG) ${\mathcal{G}}^{\prime}=({\mathcal{N}}^{\prime},{\mathcal{E}}^{\prime})$ of ${\mathcal{G}}$ is a subgraph of ${\mathcal{G}}$ (${\mathcal{N}}^{\prime}\subseteq{\mathcal{N}},{\mathcal{E}}^{\prime}\subseteq{\mathcal{E}}$) consisting of directed edges and with no cycles.
+Let the nodes of ${\mathcal{G}}^{\prime}$ be topologically ordered as ${\mathcal{N}}^{\prime}=[v_{1},v_{2}\dots,v_{l}]$.
+Let the $i^{th}$ ($i\in[1,l]$) nodes of ${\mathcal{N}}^{\prime}$ from $v_{1}$ and backwards from $v_{l}$ be ${\mathcal{N}}^{\prime}[i]\coloneqq v_{i}$ and ${\mathcal{N}}^{\prime}[-i]\coloneqq v_{l-i+1}$ respectively.
+Nodes ${\mathcal{N}}^{\prime}_{source}$ with no incoming edges are source nodes of ${\mathcal{G}}^{\prime}$.
+Nodes ${\mathcal{N}}^{\prime}_{sink}$ with no outgoing edges are sink nodes of ${\mathcal{G}}^{\prime}$.
+DAGs generalize paths that are typically used to define multi-hop knowledge comprehension queries (Yang et al., 2018; Welbl et al., 2018).
+Hence, we define multi-hop knowledge comprehension queries with DAGs in Definition 2.1.
 Definition 2.1.
-(Path in a Knowledge Graph).
-A path $\Pi=[v_{1},v_{2}\dots,v_{l}]$ is an ordered collection of nodes in a given knowledge graph ${\mathcal{G}}$, where $l>1$, such that $\forall i\in[l-1],v_{i}\in{\mathcal{N}}$, $(v_{i},v_{i+1})\in{\mathcal{E}}$, and $\forall j\in[l],i\neq j\implies v_{i}\neq v_{j}$.
-$\Pi_{H}\coloneqq v_{1}$ and $\Pi_{T}\coloneqq v_{l}$ are the head and tail nodes respectively of $\Pi$.
-Let the $i^{th}$ ($i\in[1,l]$) nodes of $\Pi$ from $\Pi_{H}$ and backwards from $\Pi_{T}$ be $\Pi[i]\coloneqq v_{i}$ and $\Pi[-i]\coloneqq v_{l-i+1}$ respectively.
-Definition 3.2 describes a multi-hop reasoning problem, derived from a given knowledge graph ${\mathcal{G}}$.
-As ${\mathcal{G}}$ naturally encodes several multi-hop problems, we use it to form the specification for a target language model ${\mathcal{L}}$, similar to prior works such as (Ho et al., 2020; Jiang et al., 2023b).
+(Multi-hop queries).
+Consider a DAG ${\mathcal{G}}^{\prime}$ in ${\mathcal{G}}$ with nodes ${\mathcal{N}}^{\prime}$.
+A multi-hop query ${\mathcal{Q}}$ for ${\mathcal{N}}^{\prime}$ is to identify the sink nodes ${\mathcal{N}}^{\prime}_{sink}$, given aliases of source nodes ${\mathcal{N}}^{\prime}_{source}$ and aliases of all edges in DAG ${\mathcal{G}}^{\prime}$.
+To avoid ambiguity regarding correct answer, we enforce ${\mathcal{N}}^{\prime}_{sink}=\{{\mathcal{N}}^{\prime}[-1]\}$ in DAGs for multi-hop knowledge comprehension queries.
+The ${\mathcal{G}}$ naturally encodes several multi-hop queries, that form the specification distributions for a language model ${\mathcal{L}}$.
+To aid reasoning and reduce dependence on parametric knowledge, prompts include alongside relevant textual context needed to identify the intermediate and final nodes.
+The context may contain natural noise, such as distractor texts or jumbled information, may occur in practical settings.
+We model such noise in our prompt distributions as described next.
+2.1.1 Natural noise in prompts
+Distractors.
+Prior works (Shi et al., 2023) on analyzing reasoning in Language Models (LMs) have shown the negative influence of irrelevant information (distractor) in prompts on LM performance.
+Hence, we include distractor texts in prompt contexts and specify that model response should not use the distractor information.
+Contexts of nodes $\tilde{v}$ adjacent to any node ${\mathcal{N}}[i]$ ($i\in[1,l]$) in ${\mathcal{N}}$ that are not the sink node or its predecessors, such that the aliases of the relation $({\mathcal{N}}[i],\tilde{v})$ are the same as those of $({\mathcal{N}}[i],{\mathcal{N}}[j])$, $j>i$ in the DAG, can serve as effective distractors for ${\mathcal{Q}}$ (Definition 2.2).
+This is because, at any intermediate step, the LM ${\mathcal{L}}$ can pick $\tilde{v}$ as the response, which can deviate ${\mathcal{L}}$’s reasoning from ${\mathcal{N}}$.
+Nodes adjacent to ${\mathcal{N}}[-1]$ and its predecessors in ${\mathcal{N}}$ are not distractors.
+For the former, the LM must have already reached the final answer before reaching its adjacent nodes, hence answering ${\mathcal{Q}}$.
+In the latter, adjacent nodes following same relation are valid answers, not distractors.
 Definition 2.2.
-(Multi-hop reasoning problem from ${\mathcal{G}}$).
-Consider any path $\Pi$ (Definition 3.1) of length $l$ in ${\mathcal{G}}$.
-A multi-hop reasoning problem ${\mathcal{Q}}$ for $\Pi$ is identifying the tail node $\Pi_{T}$, given an alias of its head node $\Pi_{H}$ and aliases of all edges from $\Pi_{H}$ to $\Pi_{T}$ in $\Pi$.
-Let $v_{\mathcal{A}}$ denote the corresponding aliases of node $v$ in ${\mathcal{G}}$.
-Let $\mathcal{D}$ be a function that samples a random alias from the given set of aliases.
-${\mathcal{Q}}\coloneqq\mathcal{D}(\Pi_{H,\mathcal{A}})\xrightarrow[]{\mathcal{D}((\Pi_{H},\Pi[2])_{\mathcal{A}})}\dots\xrightarrow[]{\mathcal{D}((\Pi[-2],\Pi_{T})_{\mathcal{A}})}?$
-${\mathcal{Q}}$ thus involves $l-1$ reasoning steps to get to the final answer, where each step requires correctly identifying intermediate nodes of $\Pi$.
-Note, however, that correctness of intermediate reasoning steps is generally not evaluated, and accuracy is defined for the final response (Rajpurkar et al., 2016).
-To aid ${\mathcal{L}}$ in correctly answering a multi-hop reasoning query ${\mathcal{Q}}$ and reduce hallucination (Dhuliawala et al., 2023), we provide relevant textual information needed to identify the intermediate and final nodes in the prompt.
-Hence, the overall task of answering ${\mathcal{Q}}$ involves information extraction for identifying intermediate nodes and reasoning to connect the intermediate answers to reach the final answer, which we collectively call knowledge comprehension.
-Our overall property quantifies the probability of observing correct knowledge comprehension for a random multi-hop reasoning query developed from ${\mathcal{G}}$.
-We formally define the property for ${\mathcal{L}}$ as a probabilistic program over ${\mathcal{G}}$ in Algorithm 1.
+(Distractor).
+Consider DAG ${\mathcal{G}}^{\prime}$ in ${\mathcal{G}}$, with nodes ${\mathcal{N}}^{\prime}$ and ${\mathcal{N}}$ respectively.
+A distractor node $\tilde{v}\notin{\mathcal{N}}^{\prime}$ satisfies $\exists v\in{\mathcal{N}}\setminus{\mathcal{N}}^{\prime},((v,\tilde{v})\in{\mathcal{E}})\wedge(\exists v^{\prime}\in{\mathcal{N}}^{\prime},(v,\tilde{v})_{\mathcal{A}}=(v,v^{\prime})_{\mathcal{A}}\wedge v^{\prime}\neq{\mathcal{N}}^{\prime}[-1])$.
+Shuffling.
+Prior works (Chen et al., 2024) have shown that LM performance can vary with information ordering.
+Hence, we shuffle the information in provided in prompt, to specify that LM’s response should be invariant to information ordering.
+Our property quantifies the probability of observing correct knowledge comprehension for a random multi-hop reasoning prompt, with possible natural noise, developed from ${\mathcal{G}}$.
+We formally define the property for ${\mathcal{L}}$ as a probabilistic program over ${\mathcal{G}}$ in Alg 1.
 We follow the syntax of the imperative probabilistic programming language in (Sankaranarayanan et al., 2013, Figure 3).
-The language has primitives for sampling from common distributions such as Uniform ($\mathcal{U}$), Bernoulli ($Ber$), etc., and an estimateProbability(.) function that outputs the probability of a random variable attaining a certain value.
-As all the random sampling steps in Algorithm 1 can operate with any discrete distribution, we use a generic identifier $\mathcal{D}$ (line 1) for samplers of discrete distributions ($\dots$’ denotes samplers for other discrete distributions).
-[…]
-A prompt for ${\mathcal{L}}$ consists of ${\mathcal{Q}}$ and a context $\Gamma$ containing information relevant to answer ${\mathcal{Q}}$.
-$\Gamma$ is formed by concatenating ($\odot$) the contexts for all nodes in $\Pi$.
-Let $v_{\gamma}$ denote the corresponding context of node $v$ in ${\mathcal{G}}$.
-Prior works (Shi et al., 2023) on analyzing reasoning in LMs have shown the negative influence of irrelevant information (distractor) in prompts on the performance of LMs, which is not ideally expected.
-Hence, we include distractor texts in $\Gamma$ and specify that the correct response should not be based on the distractor information.
-The contexts of nodes $\tilde{v}$ adjacent to any node $\Pi[i]$ ($i\in[1,l-2]$) on $\Pi$, such that the relation of $(\Pi[i],\tilde{v})$ is the same as that of $(\Pi[i],\Pi[i+1])$, can serve as effective distractors for ${\mathcal{Q}}$ (Definition 3.3).
-This is because, at any intermediate step, the model can pick $\tilde{v}$ as the response, which can deviate ${\mathcal{L}}$’s reasoning from $\Pi$.
-Nodes adjacent to $\Pi[-1]$ and $\Pi[-2]$ are not distractors.
-For the former, the model must have already reached the final answer before getting to its adjacent nodes, hence, answering ${\mathcal{Q}}$.
-In the latter, adjacent nodes following same relation are valid correct answers and not distractors.
-We denote distractor text in $\Gamma$ as the context of randomly sampled nodes from a distribution $\mathcal{D}$ over all distractor nodes of $\Pi$ in ${\mathcal{N}}$.
-We demonstrate the effects of using distractor text on the performance of SOTA LLMs in Section 4.
-Input: ${\mathcal{L}},{\mathcal{G}},v_{1},\rho$
+The language has primitives for sampling from common distributions like Uniform ($\mathcal{U}$), Bernoulli ($Ber$), etc., and estimateProb(.) function that outputs the probability of a random variable attaining a certain value.
+We use a generic identifier $\mathcal{D}$ (line 1) for samplers of discrete distributions ($\dots$’ denotes samplers for other discrete distributions).
+We use any(.) function to denote that at least $1$ of its inputs is true.
+A prompt for ${\mathcal{L}}$ consists of ${\mathcal{Q}}$ and a context $\Gamma$ containing information relevant to answer ${\mathcal{Q}}$ (Alg 1, line 6).
+${\mathcal{Q}}$ is developed from a randomly sampled DAG ${\mathcal{G}}^{\prime}$, given by the function sampleDAG.
+We elaborate on sampleDAG() in §2.1.2.
+${\mathcal{G}}^{\prime}$ is transformed into a natural language query with a randomly-selected template from a set of prespecified templates, $\tau$.
+The template samples random aliases of nodes in ${\mathcal{N}}^{\prime}_{source}$ and relations in ${\mathcal{G}}^{\prime}$ to form a natural language query for ${\mathcal{N}}^{\prime}_{sink}$.
+$\Gamma$ is formed by concatenating ($\odot$) contexts for all nodes in ${\mathcal{N}}^{\prime}$, followed by optional shuffling.
+We denote distractor text in $\Gamma$ as context of randomly sampled nodes from a distribution $\mathcal{D}$ over all distractor nodes of ${\mathcal{N}}^{\prime}$ in ${\mathcal{N}}$.
+Our final certificate (line 7) quantifies the probability that ${\mathcal{L}}$ generates any alias of ${\mathcal{N}}^{\prime}_{sink}$, which is correct answer.
+The certificate depends on the different sampling steps and is specific to ${\mathcal{G}}$.
+Input: ${\mathcal{L}},{\mathcal{G}},\tau,\texttt{args}$
 Output: $p$
 1:  $\mathcal{D}\coloneqq\mathcal{U}\mid Ber\mid...$
-2:  $\Pi\coloneqq[v_{1},v_{2}\coloneqq(\mathcal{D}([v^{\prime}\mid(v_{1},v^{\prime})\in{\mathcal{E}}])),\dots,v_{\mathcal{D}(\{2,\dots,\rho\})}\coloneqq(\mathcal{D}([v^{\prime}\mid(v_{k},v^{\prime})\in{\mathcal{E}}]))]$
-3:  ${\mathcal{Q}}\coloneqq\mathcal{D}(\Pi[0]_{\mathcal{A}})\xrightarrow[]{\mathcal{D}(\Pi[0],\Pi[1])_{\mathcal{A}}}\dots\xrightarrow[]{\mathcal{D}(\Pi[-2],\Pi[-1])_{\mathcal{A}}}\textbf{?}$
-4:  $\Gamma\coloneqq\texttt{shuffle}([\Pi[0]_{\gamma},\dots,\Pi[-1]_{\gamma},(\mathcal{D}({\mathcal{N}}))_{\gamma},\dots])$
-5:  $\mathcal{P}\coloneqq\Gamma\odot{\mathcal{Q}}$
-6:  $p\coloneqq\texttt{estimateProbability}({\mathcal{L}}(\mathcal{P})==\texttt{any}(\Pi[-1]_{\mathcal{A}}))$
-Algorithm 1 Knowledge comprehension specification
+2:  ${\mathcal{G}}^{\prime}\coloneqq\texttt{sampleDAG}({\mathcal{G}},\texttt{args})$
+3:  ${\mathcal{N}}^{\prime}\coloneqq\texttt{topologicalOrder}({\mathcal{G}}^{\prime})$
+4:  ${\mathcal{Q}}\coloneqq\mathcal{D}(\tau)({\mathcal{G}}^{\prime},{\mathcal{N}}^{\prime})$
+5:  $\Gamma\coloneqq\texttt{shuffle}([{\mathcal{N}}^{\prime}[0]_{\gamma},\dots,{\mathcal{N}}^{\prime}[-1]_{\gamma},(\mathcal{D}({\mathcal{N}}))_{\gamma},\dots])$
+6:  $\mathcal{P}\coloneqq\Gamma\odot{\mathcal{Q}}$
+7:  $p\coloneqq\texttt{estimateProb}(\texttt{any}({\mathcal{L}}(\mathcal{P})=={\mathcal{N}}^{\prime}[-1]_{\mathcal{A}}))$
+Algorithm 1 General certification
+Input: ${\mathcal{G}},\rho$; Output: ${\mathcal{G}}^{\prime}$
+1:  $l\coloneqq\mathcal{D}(\{2,\dots,\rho\})$
+2:  $\mathcal{B}\coloneqq\texttt{boundedDAGs}({\mathcal{G}},{\mathcal{N}}_{source},l)$
+3:  ${\mathcal{G}}^{\prime}\coloneqq\mathcal{D}(\mathcal{B})$
+Algorithm 2 Entity centric sampleDAG
+Input: ${\mathcal{G}},{\mathcal{G}}_{ref}$; Output: ${\mathcal{G}}^{\prime}$
+1:  $\mathcal{I}\coloneqq\texttt{generateIsomorphisms}({\mathcal{G}},{\mathcal{G}}_{ref})$
+2:  ${\mathcal{G}}^{\prime}\coloneqq\mathcal{D}(\mathcal{I})$
+Algorithm 3 Relations centric sampleDAG
+2.1.2 Local specifications
+We define individual knowledge comprehension specifications on queries over DAGs that share a common characteristic, thus forming local specifications (Seshia et al., 2018).
+Local specifications are commonly defined for and used in neural network verification (Singh et al., 2019; Baluta et al., 2021), due to their tractability to verify and simplicity to interpret.
+Although there can be various common characteristics to specify DAGs, we define two kinds of local specifications with different common characteristics.
+sampleDAG() randomly samples DAGs from given ${\mathcal{G}}$ having the common characteristics.
+[…]
+Relations centric.
+Question-answering datasets are commonly formed with question templates (Cui et al., 2016) each having fixed relations (relation-centric) between variable entities of specific types.
+For example, “Which drug is a treatment for \${disease}?”, where \${disease} can be substituted with any curable disease to form individual questions.
+We define specifications by fixing the structure and relations of DAG and randomly sampling nodes to obtain the final DAG over which (relation-centric) queries are formed.
+Alg 3 defines the sampling procedure for ${\mathcal{G}}^{\prime}$ to obtain relation-centric DAGs from reference ${\mathcal{G}}_{ref}$.
 Definition 2.3.
-(Distractor node).
-A distractor node $\tilde{v}$ for a path $\Pi=[v_{1},v_{2},\dots,v_{l}]$ of ${\mathcal{G}}$ is such that $\forall i\in[1,l],\tilde{v}\neq v_{i}$, and $\exists j\in[1,l-2],[(v_{j},\tilde{v})\in{\mathcal{E}}]\wedge[(v_{j},\tilde{v})_{\mathcal{A}}=(v_{j},v_{j+1})_{\mathcal{A}}]$.
-Prior works such as (Chen et al., 2024) have shown that LLM performance can vary with information ordering.
-Hence, we shuffle the information in $\Gamma$ (line 4) to specify that the model’s response should be invariant to the ordering of information.
-Our final specification in line 6 is the probability that ${\mathcal{L}}$ generates any alias of the last node of the path, which is the correct answer to ${\mathcal{Q}}$.
-The specification depends on the choices for the different distributions used at various sampling steps, ${\mathcal{G}}$, $v_{1}$, and $\rho$.
-It leads to certificates for the behavior of ${\mathcal{L}}$ on a given subgraph of ${\mathcal{G}}$ and paths of length at most $\rho$.
+(Isomorphic DAGs) Let finite DAGs ${\mathcal{G}}_{1}=({\mathcal{N}}_{1},{\mathcal{E}}_{1})$ and ${\mathcal{G}}_{2}=({\mathcal{N}}_{2},{\mathcal{E}}_{2})$ be isomorphic when there exists a bijection $\phi$ between the nodes of ${\mathcal{G}}_{1}$ and ${\mathcal{G}}_{2}$ such that the following conditions hold.
+(1) $\forall v_{1},v_{2}\in{\mathcal{N}}_{1},(v_{1},v_{2})\in{\mathcal{E}}_{1}\Leftrightarrow(\phi(v_{1}),\phi(v_{2}))\in{\mathcal{E}}_{2}$; (2) $\forall(v_{1},v_{2})\in{\mathcal{E}}_{1},(v_{1},v_{2})_{\mathcal{A}}=(\phi(v_{1}),\phi(v_{2}))_{\mathcal{A}}$.
+generateIsomorphisms() gives all isomorphisms of ${\mathcal{G}}_{ref}$ from ${\mathcal{G}}$.
 2.2 Certification method
-Our algorithm certifies the target LLM ${\mathcal{L}}$ by computing an interval $[p_{l},p_{u}]$ containing the value of the probability $p$ (Algorithm 1, line 6) for a given pivot node $v_{1}$ in ${\mathcal{G}}$ with high confidence.
-We model $p$ as the probability of setting the underlying boolean random variable $\mathcal{R}=({\mathcal{L}}(\mathcal{P})==\texttt{any}(\Pi[-1]_{\mathcal{A}}))$ to true (success).
+Our algorithm certifies the target LLM ${\mathcal{L}}$ by computing an interval $[p_{l},p_{u}]$ bounding probability $p$ (Alg 1, line 7) for a given specification distribution over ${\mathcal{G}}$ with high confidence.
+We model $p$ as the probability of setting the underlying boolean random variable $\mathcal{R}\triangleq({\texttt{any}({\mathcal{L}}(\mathcal{P})=={\mathcal{N}}^{\prime}[-1]_{\mathcal{A}})})$ to true (success).
 Thus, $\mathcal{R}\sim Ber(p)$.
-Exactly determining $p$ would require enumerating over all possible $\mathcal{P}$ which can be developed from any path from a subgraph of ${\mathcal{G}}$ with any random aliases, resulting in an infeasible number of possible prompts, as shown in Appendix B.6.
+Exactly determining $p$ would require enumerating over all possible $\mathcal{P}$ which can be developed from any DAG of ${\mathcal{G}}$ with any random aliases, resulting in an infeasible number of possible prompts, as shown in Appendix D.8.
 Moreover, we want our method to generalize to closed-source LLMs as well, where the internal structures of the models are unknown.
@@
 Confidence is defined as the probability of the true value $p$ being within the bounds, i.e., $Pr[p\in[p_{l},p_{u}]]$.
-To establish formal guarantees, we want our estimation procedure to be such that the actual confidence is at least the user-specified confidence level, $1-\delta$ (i.e., $Pr[p_{l}\leq p\leq p_{u}]\geq 1-\delta$), where $\delta$ is a small positive constant.
-Hence we use conservative method of the Clopper-Pearson confidence intervals (Clopper & Pearson, 1934; Brown et al., 2001; Kurz et al., 2014; Collins, 2010), which is known to produce intervals that are guaranteed to have high confidence.
-To compute the high-confidence bounds on $p$, we make $n$ independent and identically distributed observations of $\mathcal{R}$, out of which we obtain $0\leq k\leq n$ successes.
-With the $n$ observations and $1-\delta$ confidence, we use Clopper-Pearson confidence intervals to bound $p$.
+To establish formal guarantees, we want our estimation procedure to be such that the actual confidence is at least the user-specified confidence level, $1-\delta$ (i.e., $Pr[p_{l}\leq p\leq p_{u}]\geq 1-\delta$), where $\delta>0$ is a small constant.
+Hence, we use Clopper-Pearson (CP) confidence intervals (Clopper & Pearson, 1934; Kurz et al., 2014), which is a conservative method known to produce guaranteed high confidence intervals.
+We show empirical evidence of the bounds’ conservativeness in Appendix C.
+To compute high-confidence bounds on $p$, we make $n$ independent and identically distributed observations of $\mathcal{R}$, in which we obtain $k$ successes, $k\in[0,n]$.
+We generate CP confidence intervals with the $n$ observations to bound $p$ with $1-\delta$ confidence.
 3 Experiments
-4.1 Experimental setup
-We certify the following open-source, instruction finetuned (Wei et al., 2022a) models — Llama-3-instruct 8B model (Dubey et al., 2024), Mistral 7B-Instruct-v0.2 (Jiang et al., 2023a), Phi-3 3B and 14B parameter models (Abdin et al., 2024).
-We also certify 4-bit and 8-bit quantized versions of the open-source models to study the effects of quantization on a model’s knowledge comprehension capabilities.
-Among the closed-source models with API access, we certify Gemini-1.5 (Gemini Team, 2024) Flash-001 and Pro-002 models and GPT-4o-0827 (OpenAI, 2024).
-We use Wikidata5m (Wang et al., 2021) as our knowledge graph after preprocessing (check Appendix B.1.1 for details).
-To generate challenging and diverse specifications, we sample 50 pivot nodes from two populations: the top 2000 nodes by out-degree in the global graph, and nodes whose subgraph within radius $\rho$ contains at least 2000 vertices.
-This strategy ensures specifications rooted around any of the pivot nodes have a large number of paths, making enumerative certification (where all possibilities are tested for satisfaction of the specification) impractical.
-We set the maximum path length parameter as $\rho=5$, as we empirically observe that longer paths could result in queries that are very unrelated to the head node of the path.
-As our certificates are over all paths with lengths at most $\rho$ in a given subgraph, we equally prioritize the different possible path lengths in $[1,\rho]$, even though paths with longer lengths can be fewer in number than those with shorter lengths.
-Hence, we define our sampler from our distribution over paths (Algorithm 1, line 2) which first selects a path length from $[1,\rho]$ uniformly randomly and then selects a path having the chosen path length also uniformly randomly.
-A query is constructed from the selected path by uniformly selecting any aliases for the nodes in the path (Algorithm 1, line 3).
-We use uniform distributions so as not to bias the results by prioritizing some elements of the underlying sample spaces.
-Note, however, the distributions can be modified according to the specific certification usecases.
-We study the variation in the performance of LLMs with varying path lengths within $[1,\rho]$ in Appendix A.
-A prompt is constructed for a selected path and query by prepending a context containing the relevant information to answer the query.
-The design of the context results in $3$ kinds of specifications, which we describe in Section 4.2.
-We provide some few-shot examples in the context as well, to explain the task to the models.
-We provide an ablation study on the effects of varying the number of few-shot examples on the performance of the target LLM in Appendix A.
-We formulate our prompts as multiple choice questions to facilitate evaluation of the LLM responses.
-The distribution of prompts is primarily determined by the query generation and answer option selection processes, with occasional context adjustments to accommodate model-specific context window constraints.
-Queries are derived from paths with uniformly distributed lengths, ensuring a balanced representation of reasoning complexity.
-Answer options are sampled non-uniformly, prioritizing distractors (in the distractor setting), followed by entities from the query path, and finally random entities related to path nodes.
-This approach balances query complexity and creates challenging answer options.
-We detail the complete prompt construction process in Appendix B.3.
-For evaluation of LLM responses, we use a string matching procedure.
-The detailed design of our evaluator is presented in Appendix B.7.
-QuaCer-C generates certificates with confidence $1-\delta=95\%$ and number of samples, $n=250$ samples.
-We conduct our experiments for open-source LLMs on 2 A100 GPUs.
-4.2 Different kinds of specifications
-We study the certificates for $3$ different kinds of specifications that arise from variations in the construction of context in the prompts to LLMs (Algorithm 1, line 4) — with context shuffling and distractor context (Shuffle Distractor), with context shuffling and without distractor context (Shuffle), and without context shuffling and without distractor context (Vanilla).
-These settings enable us to study the effects of these operations on the knowledge provided in LLM prompts.
-When distractor context is provided, it is only for $1$ distractor node, so as to fit the relevant context for the nodes on the considered path within the context windows of the target LLMs.
-We hypothesize that distractors to nodes later in the path, closer to the tail node which consists of the final answer, would be more challenging for the LLM due to their proximity and similarity to the answer node.
-Our distractor node sampler (Appendix B.3, Algorithm 4) thus employs a weighted sampling approach to prioritize distractor nodes closer to the path’s tail.
-We provide an ablation study on the effects of varying the distribution of distractor nodes in Appendix A.
-Table 1: Certification results for different LLMs
-Model Specification kind Precision Avg. lower bound Avg. upper bound Avg. accuracy
-Gemini-1.5-Pro Vanilla - $0.72\pm 0.06$ $0.83\pm 0.05$ $0.78\pm 0.06$
-Shuffle - $0.71\pm 0.06$ $0.82\pm 0.05$ $0.77\pm 0.06$
-Shuffle Distractor - $0.64\pm 0.09$ $0.75\pm 0.08$ $0.70\pm 0.09$
-GPT-4o Vanilla - $0.70\pm 0.06$ $0.81\pm 0.06$ $0.76\pm 0.06$
-Shuffle - $0.69\pm 0.06$ $0.80\pm 0.06$ $0.75\pm 0.06$
-Shuffle Distractor - $0.62\pm 0.09$ $0.74\pm 0.08$ $0.68\pm 0.09$
-Gemini-1.5-Flash Vanilla - $0.46\pm 0.06$ $0.58\pm 0.06$ $0.52\pm 0.06$
-Shuffle - $0.45\pm 0.06$ $0.57\pm 0.06$ $0.51\pm 0.06$
-Shuffle Distractor - $0.42\pm 0.10$ $0.55\pm 0.10$ $0.48\pm 0.10$
-Llama (8B) Vanilla fp16 $0.31\pm 0.09$ $0.43\pm 0.10$ $0.36\pm 0.10$
-8bit $0.30\pm 0.05$ $0.42\pm 0.06$ $0.36\pm 0.06$
-4bit $0.27\pm 0.05$ $0.39\pm 0.05$ $0.33\pm 0.05$
-Shuffle fp16 $0.31\pm 0.05$ $0.44\pm 0.06$ $0.37\pm 0.06$
-8bit $0.31\pm 0.06$ $0.44\pm 0.06$ $0.37\pm 0.06$
-4bit $0.28\pm 0.07$ $0.40\pm 0.07$ $0.34\pm 0.07$
-Shuffle Distractor fp16 $0.30\pm 0.10$ $0.42\pm 0.11$ $0.36\pm 0.11$
-8bit $0.30\pm 0.06$ $0.42\pm 0.06$ $0.35\pm 0.06$
-4bit $0.25\pm 0.09$ $0.36\pm 0.09$ $0.30\pm 0.09$
-Mistral (7B) Vanilla fp16 $0.33\pm 0.05$ $0.45\pm 0.05$ $0.39\pm 0.05$
-8bit $0.32\pm 0.05$ $0.44\pm 0.06$ $0.38\pm 0.05$
-4bit $0.31\pm 0.06$ $0.43\pm 0.06$ $0.37\pm 0.06$
-Shuffle fp16 $0.34\pm 0.05$ $0.46\pm 0.06$ $0.40\pm 0.05$
-8bit $0.34\pm 0.06$ $0.46\pm 0.06$ $0.40\pm 0.06$
-4bit $0.32\pm 0.05$ $0.44\pm 0.06$ $0.38\pm 0.06$
-Shuffle Distractor fp16 $0.33\pm 0.05$ $0.45\pm 0.05$ $0.39\pm 0.05$
-8bit $0.33\pm 0.11$ $0.46\pm 0.12$ $0.39\pm 0.12$
-4bit $0.28\pm 0.11$ $0.39\pm 0.12$ $0.33\pm 0.11$
-Phi-3 (14B) Vanilla fp16 $0.35\pm 0.05$ $0.47\pm 0.05$ $0.41\pm 0.05$
-8bit $0.35\pm 0.05$ $0.47\pm 0.05$ $0.41\pm 0.05$
-4bit $0.33\pm 0.04$ $0.45\pm 0.05$ $0.39\pm 0.05$
-Shuffle fp16 $0.35\pm 0.04$ $0.48\pm 0.04$ $0.41\pm 0.04$
-8bit $0.34\pm 0.06$ $0.47\pm 0.06$ $0.40\pm 0.06$
-4bit $0.33\pm 0.04$ $0.46\pm 0.05$ $0.39\pm 0.05$
-Shuffle Distractor fp16 $0.33\pm 0.11$ $0.45\pm 0.11$ $0.38\pm 0.11$
-8bit $0.31\pm 0.08$ $0.43\pm 0.09$ $0.37\pm 0.08$
-4bit $0.30\pm 0.08$ $0.42\pm 0.09$ $0.36\pm 0.09$
-Phi-3 (3B) Vanilla fp16 $0.34\pm 0.05$ $0.46\pm 0.06$ $0.40\pm 0.05$
-8bit $0.32\pm 0.05$ $0.44\pm 0.06$ $0.38\pm 0.06$
-4bit $0.32\pm 0.05$ $0.44\pm 0.06$ $0.38\pm 0.06$
-Shuffle fp16 $0.34\pm 0.04$ $0.47\pm 0.05$ $0.40\pm 0.05$
-8bit $0.32\pm 0.04$ $0.44\pm 0.05$ $0.38\pm 0.05$
-4bit $0.32\pm 0.05$ $0.44\pm 0.05$ $0.38\pm 0.05$
-Shuffle Distractor fp16 $0.32\pm 0.10$ $0.45\pm 0.10$ $0.38\pm 0.10$
-8bit $0.31\pm 0.09$ $0.43\pm 0.10$ $0.37\pm 0.10$
-4bit $0.29\pm 0.10$ $0.41\pm 0.10$ $0.35\pm 0.10$
+Our experiments were run on 2 A100 GPUs with 40GB VRAM each.
+We certify the following open-source models — Llama-3.2-instruct 1B, 3B, 11B (Dubey et al., 2024), Phi-3 3B and 14B (abdin2024Phi-3technicalreporthighly) and their 4-bit quantizations to study the effects of quantization on knowledge comprehension.
+Among closed-source models with API access, we certify Gemini-1.5 (Gemini Team, 2024) Flash-001 and Pro-002 and GPT-4o-0827 (OpenAI, 2024).
+We use PrimeKG (Chandak et al., 2023) and Wikidata5m (Wang et al., 2021b) as our knowledge graphs (KGs) (details in Appendix D.3.1 and Appendix D.2.1 respectively).
+PrimeKG is a KG for precision medicine, while Wikidata5m is for general knowledge from Wikipedia.
+These practical KGs are chosen only for illustration, and our framework generalizes to other KGs.
+We show relation-centric and entity-centric specifications for PrimeKG and Wikidata5,m respectively.
+PrimeKG.
+We manually define 9 relation-centric specifications based on common clinical reasoning patterns.
+All specifications are shown in §D.4.2.
+Each specification is defined by: 1.
+A reference DAG (${\mathcal{G}}_{ref}$); 2.
+A set of natural language templates (5-7 per specification); 3.
+Constraints on entity types for isomorphic DAG sampling.
+An example specification is the “drug-disease interaction” specification.
+It consists of ${\mathcal{G}}_{ref}$ which is a 3-node DAG (shown on the right) with drug→disease edges.
+An example template to form natural language queries from ${\mathcal{G}}_{ref}$ is “Which drug treats disease B but is contraindicated for disease A?”.
+Wikidata5m.
+We created 50 entity-centric specifications using Wikidata5m, sampling linear paths (DAGs) up to length $\rho=5$ to maintain query relevance.
+The 50 starting nodes were selected from high out-degree (top 2k) or high-density neighborhood ($>$2k nodes within radius $\rho$) populations the specifications have a large number of paths, making enumeration infeasible.
+Note that these specifications are only for demonstration and LLMCert-C is not limited to them.
+To avoid bias towards the naturally more frequent shorter paths, we first sample a desired path length uniformly from $[2,\rho]$.
+The actual path is then generated from the starting node according to this length (Alg.
+2).
+We analyze the impact of varying query path lengths on LLM performance in Appendix E.
+Note that the framework is adaptable to other distributions as needed by the specific certification usecase.
+Specifications with varying complexities.
+We certify $2$ kinds of specifications with different kinds of contexts in LLM prompts (Alg 1, line 4) — with context shuffling and distractor context (Distractor) and without such noise (Vanilla).
+This enables us to study the effects of natural noise on LLM performance.
+Leveraging the linear structure of Wikidata5m DAGs, we use weighted sampling (Alg.
+6) to select distractors closer to the answer node, hypothesizing this increases difficulty.
+Since PrimeKG lacks such unique linearizations, its distractor nodes are sampled uniformly.
+We provide an ablation study on the effects of varying the weighted distractor sampling in Appendix E.
+Prompts give multiple answer options with $1$ correct, to simply evaluate LLM responses using string matching (Appendix D.9).
+For answer options in Distractor setting, we prioritize distractor nodes, followed by nodes from the query path, and finally, random nodes related to path nodes.
+Vanilla setting is similar, without distractors.
+All further details of prompt construction are in Appendix D.5.
+Baseline.
+We compare LLMCert-C’s results with a benchmarking baseline.
+It measures the target LLM’s accuracy on a static dataset.
+We generate 50 static queries from each specification by uniformly sampling DAGs.
+Queries use a fixed alias of entities and relations, with no shuffling or distractors in contexts to remove natural noise.
+Table 1: Average certification bounds for specifications over Wikidata5m and PrimeKG KGs for different LLMs. Bounds are given as [lower bound, upper bound].
+Model Quantization Wikidata PrimeKG
+Baseline Vanilla Distractor Baseline Vanilla Distractor
+Gemini-1.5-Pro - $0.87$ [0.80, 0.89] [0.64, 0.75] $1.00$ [0.97, 1.00] [0.92, 0.98]
+GPT-4o - $0.88$ [0.80, 0.89] [0.62, 0.74] $0.99$ [0.97, 1.00] [0.85, 0.95]
+Gemini-1.5-Flash - $0.81$ [0.72, 0.83] [0.43, 0.56] $0.99$ [0.96, 0.99] [0.89, 0.96]
+Phi-3 (14B) fp16 $0.63$ [0.54, 0.66] [0.33, 0.45] $0.98$ [0.93, 0.98] [0.75, 0.85]
+4bit $0.63$ [0.52, 0.65] [0.30, 0.42] $0.95$ [0.92, 0.97] [0.71, 0.81]
+Llama (11B) fp16 $0.61$ [0.50, 0.63] [0.33, 0.45] $0.97$ [0.91, 0.97] [0.84, 0.92]
+4bit $0.54$ [0.45, 0.57] [0.29, 0.41] $0.94$ [0.88, 0.95] [0.79, 0.88]
+Phi-3 (3B) fp16 $0.58$ [0.50, 0.61] [0.32, 0.45] $0.92$ [0.85, 0.92] [0.70, 0.81]
+4bit $0.57$ [0.48, 0.60] [0.29, 0.41] $0.89$ [0.83, 0.91] [0.74, 0.83]
+Llama (3B) fp16 $0.51$ [0.43, 0.55] [0.27, 0.39] $0.87$ [0.80, 0.89] [0.70, 0.80]
+4bit $0.46$ [0.39, 0.51] [0.24, 0.35] $0.84$ [0.73, 0.83] [0.63, 0.73]
+Llama (1B) fp16 $0.38$ [0.30, 0.41] [0.22, 0.33] $0.48$ [0.43, 0.56] [0.39, 0.52]
+4bit $0.34$ [0.27, 0.38] [0.20, 0.31] $0.38$ [0.33, 0.45] [0.28, 0.40]
 3.1 Certificates
-QuaCer-C generates certificates providing high-confidence, tight lower and upper bounds on the probability of a correct LLM response to a random prompt sampled using the prompt constructor from the given distribution of prompts in our specifications.
-We report the average value of the lower and upper bounds, over our set of specifications that QuaCer-C certifies for each LLM, in Table 1.
-We also report the average empirical probability (the ratio of correct responses to the total number of prompts, $n$, for each certificate), averaged over the test set.
-Next, we summarize the key observations from the certification results in Table 1, some of which follow trends from prior works.
-Scaling of knowledge comprehension with model size.
-We observe that the larger models such as Gemini and GPT have significantly higher bounds than those for the smaller models such as Phi-3, Mistral, Llama (Figure 3).
-Figure 3: Variation in certification bounds with models (Vanilla, fp16)
+LLMCert-C generates certificates with high confidence ($1-\delta=95\%$), tight lower and upper bounds on the probability of a correct LLM response to a random prompt sampled from the prompt distribution in our specifications.
+LLMCert-C uses $n=250$ number of samples.
+We report the average value of the lower and upper bounds over all specifications over a KG that LLMCert-C certifies for each LLM, in Table 1.
+Each certificate takes 10-20 minutes for generation, where the main bottleneck is each LLM inference.
+We plan to optimize this in future work.
+We also report the average baseline result for each static dataset from the specifications.
+Appendix E presents detailed ablation studies for the impact of the various certification hyperparameters.
+Next, we summarize the key observations from Table 1.
+Figure 2: Samples from PrimeKG’s Distractor specifications.
+Prompt context not shown for brevity.
+Wrong/correct responses are colored red/green respectively.
+Samples are consistent with certificates, where Phi-3 (14B) has lower bounds than GPT-4o and Gemini-Pro.
+Figure 3: Negative effects of distractors on LLM performance for answering queries from same DAGs (correct and wrong answers in green and red respectively).
+Variation in knowledge comprehension with models.
+We observe that larger models, such as Gemini and GPT, have significantly higher bounds than those for the smaller models, such as Phi-3, Llama, for specifications over both KGs.
 The lower bounds of the larger models are higher than the upper bounds of the smaller models, suggesting a paradigm shift in knowledge comprehension capabilities (especially for Gemini-1.5-Pro and GPT-4o).
-However, as the larger models are also closed-source, we are unaware whether the enhanced knowledge comprehension capabilities could be due to specialized training or finetuning techniques applied on the models.
-We see that the smaller models have similar certification bounds.
-Interestingly, Phi3-3B, which is the smallest model we consider, is performing comparatively to its 14B counterparts and also to larger Mistral and Llama models.
+However, as the larger models are also closed-source, we are unaware whether the enhanced knowledge comprehension capabilities could be due to specialized training or fine-tuning techniques applied on the models.
+From the certification bounds for PrimeKG, we can establish performance hierarchies with high confidence across the evaluated models.
+Specifically, when we observe higher lower bound of a model than the upper bound of another model (e.g., Gemini-v1.5-Pro vs Llama 11B for all settings), we conclude the superior performance of the former in the worst case with high confidence.
+Effects of different kinds of specifications.
+Our results for the different kinds of specifications indicate that Vanilla specifications are easier and lead to higher certification bounds.
+Distractor specifications are challenging for all models, leading to upper bounds that are consistently lower than the lower bounds for the Vanilla setting.
+This suggests susceptibility of SOTA LLMs to natural noise in prompts.
+Even larger models such as Gemini-1.5-Pro, which show close to perfect performance in the baseline and Vanilla case (for PrimeKG only), can be incorrect with distractors, showing the nontrivial effects of noise.
+Comparison with benchmarking baseline.
+Baseline scores of all models approach or surpass average certification upper bounds, suggesting inflation of performance estimates in benchmarking.
+For example, contrary to certification bounds for PrimeKG’s distractor specifications, Phi-3 (14B) outperformed Llama (11B).
+Such findings emphasize the need for more reliable evaluation methods grounded with statistical guarantees.
+Our certificates also improve upon the baseline by preventing incorrect comparisons between models with similar performance (e.g., Phi-3 14B and Llama 11B and respectively between their 4-bit models for Vanilla specifications) by showing overlaps in their certification bounds.
 Effects of model quantization.
-We see that with increase in quantization, the model performance deteriorates on knowledge comprehension, which is contrary to prior findings such as Jin et al.
+We see that higher quantization deteriorates model performance on knowledge comprehension, contrary to prior works like Jin et al.
 (2024) that suggest that 4-bit quantization can retain the model’s knowledge and reasoning capabilities.
-The drop in the performance is not large, however.
-Effects of different kinds of specifications.
-Our results for the different kinds of specifications — Vanilla, Shuffle, and Shuffle Distractor, indicate that the Vanilla specifications are generally easier for models, leading to higher certification bounds than those for the other specifications.
-Shuffle Distractor specifications are challenging specifications for all models resulting in consistently lower certification bounds.
-However, the differences in the bounds’ values are not high across the settings, which could be because of the challenge of large and unstructured contexts from which the model need to identify relevant information in all cases.
-Quality of bounding intervals.
-Table 1 presents the average bounds from the confidence intervals generated for each certificate.
-A desirable property for the intervals, alongside their high confidence, is that they should be tight, i.e., their range should be small.
-Tighter intervals indicate precise analysis with small errors.
-We observe that the average range of the confidence intervals obtained in our experiments is less than $0.12$.
 3.2 Case studies
-Figure 4: Qualitative analysis of the samples used for certifying knowledge comprehension for the Vanilla specifications on the Wikidata5m subgraph pivoted at the node for ‘Batman Begins’ movie.
-The context provided in the prompts is not shown for brevity.
-Wrong model responses are colored red and the correct ones are colored green.
-The samples are consistent with the certification results, wherein Phi-3 (3B) has lower certification bounds than those for GPT-4o, which has lower bounds than those for Gemini-Pro.
-Figure 5: Categorizing predominant types of LLM responses. The LLMs can either do good reasoning or show failure instances caused by distractors or missed reasoning steps.
-In this section, we conduct a qualitative analysis of the certification results in Table 1.
-Firstly, we show the responses of $3$ models in Figure 4 — Phi-3 (3B), GPT-4o, and Gemini — that were obtained when certifying them for the Vanilla specification defined over a subgraph pivoted at the node for ‘Batman Begins’ movie.
-The samples are representative of the models’ certification results.
-Next, we categorize and identify the predominant kinds of model responses and present their examples.
-In particular, we frequently see the following two failure modes — distracted and missed relation.
-In the former case, the model gets deviated from the query by following the distractor context in its prompt, resulting in an incorrect answer.
-In the latter, the model skips some reasoning steps to arrive at the final, correct answer.
-In other cases of good reasoning, the model accurately follows the steps of the query and arrives at the correct answer.
-Figure 5 presents examples of the aforementioned kinds of model responses for the GPT-4o model.
+Next, we qualitatively analyze certificates for PrimeKG specifications.
+First, we show the responses of $3$ models in Figure 2 — Phi-3 (3B), GPT-4o, and Gemini-Pro, obtained when certifying them for the Distractor specification.
+The samples reflect the certification bounds, showing the performance hierarchy across the models.
+The certification bounds can be used to select the best models for critical applications.
+Next, we identify the vulnerabilities of SOTA LLMs to distractor texts in Figure 3.
+We see consistent occurrences of incorrect model responses in the presence of distractors and not otherwise.
+This demonstrates the vulnerabilities of SOTA LLMs and the risks posed by their deployment in critical applications.
+We discuss similar case studies for Wikidata5m specifications in Appendix F.
 4 Related works
 In-context learning.
-As the context windows of LLMs increase (Gemini Team, 2024; Chen et al., 2023; Dubey et al., 2024), more information can be provided to the models to improve their performance.
-Moreover, few-shot demonstrations (Brown et al., 2020) and examples from related tasks (Qu et al., 2024) can be provided in the context to inform the models about task-specific details, which can result in better performance.
-This emergent behavior (Wei et al., 2022b; Lu et al., 2024) wherein LLMs can become proficient at a task with few demonstrations provided in their prompts is in-context learning.
-We use in-context learning to enhance LLMs’ knowledge to answer the prompts and provide few-shot demonstrations for the reasoning steps expected from the models.
+As LLM context windows increase (Gemini Team, 2024; Chen et al., 2023; Dubey et al., 2024), more information can be provided in the prompts like few-shot demonstrations (Brown et al., 2020) and examples from related tasks (Qu et al., 2024).
+In-context learning is the emergent behavior (Wei et al., 2022; Lu et al., 2024) in which LLMs become proficient at a task with demonstrations in prompts.
+We use in-context learning and few-shot examples to enhance LLMs’ knowledge comprehension.
 Benchmarking LLM intelligence.
-Several benchmarks have been proposed to study the reasoning (Zhou et al., 2022; Huang & Chang, 2023; Plaat et al., 2024), arithmetic (Yuan et al., 2023; Song et al., 2024; Yang et al., 2023a), planning (Pallagani et al., 2023; Valmeekam et al., 2023; Huang et al., 2022), and question-answering (Yang et al., 2018; Ho et al., 2020; Welbl et al., 2018) capabilities of LLMs, which are integral components of human intelligence.
-These benchmarks provide empirical insights and trends into the performance of LMs.
-However, these insights are generally for static datasets and are not guaranteed to generalize.
-On the other hand, certification methods provide guarantees on, for example, the scope (defined by specifications) and confidence of its claims, as we illustrate in this work.
-6 Conclusion and future work
-In this work, we present a novel framework to formally certify LLMs for knowledge comprehension.
-We develop novel specifications that quantify the probability of correct responses over any random knowledge comprehension prompts from a distribution derived from a knowledge graph.
-Certificates for knowledge comprehension consist of high-confidence bounds on the probability of correct knowledge comprehension, thus providing a method to compare different LLMs with formal guarantees.
-Our experiments study the variations in knowledge comprehension along the axes of model size, quantization, and task difficulty and present novel trends.
-As our work is the first to certify knowledge comprehension in LLMs to the best of our knowledge, there are several possible directions for future work.
-Our framework can be integrated with knowledge graph construction methods (Ye et al., 2022), to specify and certify LLMs for comprehension and reasoning over less structured/proprietary inputs.
-The certification method can also be made more sample-efficient.
-Moreover, our framework can be adapted to domains with other kinds of graphs, such as social network graphs or medical graphs, to certify LLMs for comprehension in domains of social importance.
-Appendix A Ablations
-A.1 Few Shot Prompts
-We conduct an ablation study to examine the impact of varying the number of few-shot examples on Gemini-Flash’s performance in the vanilla task setting.
-While our default configuration uses two few-shot examples, we extend this analysis to include up to five examples.
-Interestingly, we observe no significant variation in performance across these different few-shot configurations.
-The results are presented below in  2.
-Table 2: Certification results for different LLMs with different number of few-shot examples
-Model Avg. lower bound Avg. upper bound Avg. accuracy
-Gemini-1.5-Flash Vanilla (Default) $0.46\pm 0.06$ $0.58\pm 0.06$ $0.52\pm 0.06$
-Gemini-1.5-Flash 3Shot Vanilla $0.46\pm 0.06$ $0.58\pm 0.06$ $0.52\pm 0.06$
-Gemini-1.5-Flash 4Shot Vanilla $0.46\pm 0.07$ $0.58\pm 0.07$ $0.52\pm 0.07$
-Gemini-1.5-Flash 5Shot Vanilla $0.46\pm 0.07$ $0.58\pm 0.07$ $0.52\pm 0.07$
-A.2 Distractor Distributions
-To assess the impact of distractor distribution on model performance, we implement three distinct distractor distribution strategies:
-  1. 1.
-    Tail-weighted: Linearly increasing weights towards the tail end of the path, prioritizing distractors near the answer node. This serves as our default setting.
-  2. 2.
-    Head-weighted: Linearly increasing weights towards the head of the path, emphasizing distractors near the query’s starting point.
-  3. 3.
-    Uniform: Equal probability of selecting distractors from any position along the path.
-We observe no significant differences in either of the settings. The results are presented in  3 below.
-Table 3: Certification results for Gemini-Flash with different distractor distributions
-Model Avg. lower bound Avg. upper bound Avg. accuracy
-Gemini-1.5-Flash Setting 1 (Default) $0.42\pm 0.10$ $0.55\pm 0.10$ $0.48\pm 0.10$
-Gemini-1.5-Flash Setting 2 $0.42\pm 0.11$ $0.55\pm 0.11$ $0.48\pm 0.11$
-Gemini-1.5-Flash Setting 3 $0.42\pm 0.11$ $0.55\pm 0.11$ $0.48\pm 0.11$
-A.3 Model Performances with varying Path Length
-Among our certificates, we have queries of various lengths.
-Here we study the effects on models behavior on queries with varying length by considering the number of hops they require to reason to answer the query(which is 1 less than the path length).
-To do so, we refer to the number of hops to answer a query as k where $1\leq$ k $<\rho$.
-Varying Setting: In figure  6 we show plots for various specifications for the GPT4o model.
-Figure 6: Variations in the bounds against the path lengths across various specifications.
-Varying Quantization: In figure  7 we show plots when the quantization is varied with the Llama3-8B model on the shuffle specification and their effects on performance.
-Figure 7: Variations in the bounds against the path lengths across various quantizations.
-Varying Models: In fig  8 we show plots for the shuffle specification and performance across the models(the open-source models use fp16 precision).
-Figure 8: Variations in the bounds against the path lengths across various models in the shuffle setting.
-In Figure  6, we observe that the performance across settings converges as k increases and the distractor setting is most impactful on the performance for $k=2$.
-In Figure  7, we infer that as k increases the performance of the models’ on the task converges across the different quantizations.
-We hypothesize this is due to the increasing complexity of the reasoning task.
-In Figure 8, we see that larger models (GPT-4o, Gemini-Pro) show less severe drop in performance compared to their smaller models.
-The figure shows that large models may have learnt to better apply 1-step reasoning for multiple steps when compared to their smaller counterparts.
-Appendix B Knowledge Graph and Query Generation
+Several benchmarks have been proposed to study reasoning (Huang & Chang, 2023; Plaat et al., 2024; He et al., 2024; Zha et al., 2021), arithmetic (Yuan et al., 2023; Song et al., 2024), planning (Pallagani et al., 2023; Valmeekam et al., 2023), and question-answering (Yang et al., 2018; Welbl et al., 2018; Perevalov et al., 2022) in LLMs.
+These benchmarks provide empirical insights and trends of LM performance.
+However, they use static datasets and not guaranteed to generalize.
+Certification provides guarantees on the scope (defined by specifications) and confidence of the claims, as shown in this work.
+Appendix A A note for practitioners
+Our work takes the first step towards providing formal guarantees of LLMs’ knowledge comprehension capabilities, with particular relevance for high-stakes domains like healthcare.
+While our framework enables more reliable LLM deployment through rigorous guarantees, these certificates should be viewed as probabilistic bounds rather than absolute guarantees of model reliability.
+We anticipate this work will help make LLMs more trustworthy for safety-critical applications, although continued human supervision remains essential.
+Future work can extend our framework to certify additional properties like safety and robustness, and integrate it with knowledge graph construction methods to handle less structured inputs.
+With a one-time effort to construct domain-specific knowledge graphs, practitioners can obtain reliable insights into the performance of LLMs for question-answering and knowledge comprehension tasks.
+We release our implementation to support further research on formal verification of language models.
+Appendix B Background
+B.1 Knowledge graph
+A knowledge graph ${\mathcal{G}}=({\mathcal{N}},{\mathcal{E}})$ is a collection of nodes ${\mathcal{N}}$ representing entities, interconnected by directed edges ${\mathcal{E}}$ representing their relations (Peng et al., 2023; Ji et al., 2022).
+They are commonly used in search engines to improve the accuracy of responses to user queries.
+Hence, big companies develop their own closed-source knowledge graphs.
+PrimeKG (Chandak et al., 2023) is an open-source knowledge graph of precision medicine consisting of nodes corresponding to diseases, drugs, effects, etc.
+PrimeKG describes $\sim$17k diseases with $\sim$4M relations.
+Wikidata5m (Wang et al., 2021b), another popular open-source knowledge graph consisting of 5M nodes, is a structured representation of Wikipedia.
+Each Wikidata node corresponds to a Wikipedia page, containing its abstract and a set of aliases that can synonymously refer to the node.
+Two nodes $(v_{1},v_{2})$, $v_{1},v_{2}\in{\mathcal{N}}$ are connected by a labeled edge if there is a link in the supporting document for $v_{1}$ to that for $v_{2}$.
+Edge $(v_{1},v_{2})$ is labeled by a set of aliases for the relation between $v_{1}$ and $v_{2}$.
+App.
+Figure 4 shows a subgraph of Wikidata5m.
+Figure 4: A subgraph of Wikidata5m
+B.2 Large Language Models
+Large Language Models (LLMs) are autoregressive causal language models that operate on a vocabulary $\mathcal{V}$, a set of tokens.
+LLMs takes a sequence of tokens $x_{1},...,x_{n}$ where $x_{i}\in\mathcal{V},n>0$ and outputs a probability distribution over $\mathcal{V}$ to sample the next token $x_{n+1}$.
+These models are pretrained on vast corpora (Liu et al., 2024) and have shown remarkable capabilities  (Touvron et al., 2023; Gemini Team, 2024; OpenAI, 2024).
+Numerous benchmarks (Yang et al., 2018; Rein et al., 2023; Hendrycks et al., 2021) have been developed to evaluate the performance of LLMs on tasks related to multi-step reasoning, knowledge comprehension and question answering.
+However, there remains a gap in our theoretical understanding of LLMs’ capabilities.
+Appendix C Validity of Certification bounds
+In this section, we provide empirical evidence for the validity of the certification bounds with the user-defined confidence (95%).
+We select $10$ equally-spaced values of the true probability of correct LLM response, between 0 and 1.
+For each true probability, we generate $1000$ Clopper-Pearson confidence intervals, which are used to form our certification bounds.
+We calculate the proportion of the confidence intervals that contain the true probability and compare with the confidence level.
+We show in Figure 5 that the proportion of instances where the confidence intervals contain the true probability is almost always more than $95\%$.
+Figure 5: Variation in the proportion of Clopper-Pearson confidence intervals containing the true probability of correct response
+Appendix D Knowledge Graph and Query Generation
 This section details our experimental setup for generating multi-hop reasoning queries using the Wikidata5m knowledge graph.
 We describe the structure of the knowledge graph, the process of generating random paths, formulating queries, and creating answer options including distractors.
-B.1 Knowledge Graph Structure
-Our experiments are based on the Wikidata5m knowledge graph (KG). The KG has the following key characteristics:
-  • Nodes: Each node represents an entity and is associated with a text paragraph from Wikidata5m.
-  • Edges: Edges represent relationships between entities.
-  • Text Paragraphs: The text associated with each node may contain information relevant to its connected edges.
-  • Node and Edge Aliases: Each node and each edge has a set of aliases associated with them which are just different names for them.
-This structure allows us to generate queries that require reasoning across multiple hops in the graph.
-B.1.1 Preprocessing the wikidata5m knowledge graph
+D.1 Knowledge Graph Structure
+Our experiments utilize two knowledge graphs: Wikidata5m and PrimeKG, each with distinct characteristics.
+D.2 Wikidata5m
+Wikidata5m is characterized by rich textual and semantic features:
+  • Nodes: Each node represents an entity and is accompanied by a descriptive text paragraph that provides contextual information about the entity.
+  • Edges: Relationships between entities are represented as edges.
+  • Text Paragraphs: Each node contains an associated text paragraph that that often provides at minimum information relevant to its relationships with other entities.
+  • Aliases: Each node and each edge has a set of aliases associated with them which are just different names for them.
+D.2.1 Preprocessing the wikidata5m knowledge graph
 To ensure the generation of unambiguous queries and support the certification process, we preprocess the wikidata5m dataset.
@@
     Unicode to ASCII:For consistency within our experiments, we convert all text containing Unicode characters into their respective ASCII approximations.
-B.2 Query Generation
-We utilize the Wikidata5m knowledge graph for multi-hop query generation. The query generation process involves the following steps:
-B.2.1 Random Path Generation
-We begin by selecting a pivot node $v_{0}$ in the knowledge graph ${\mathcal{G}}$.
-From this pivot, we construct a local subgraph ${\mathcal{G}}(v_{0})$ consisting of all paths $\Pi_{v_{0}}$ originating from $v$.
-This local subgraph serves as the domain for our path generation process.
-As arbitrary long paths can lose their semantic meaningfulness, we use a constraint $\rho$ to restrict the length of paths from the pivot node in the subgraph to be maximum $\rho$.
-Within ${\mathcal{G}}(v_{0})$, we generate a path $\Pi$ using a randomized depth-first search algorithm.
-The length of this path, denoted as $k_{choice}$, is sampled randomly from the set ${1,2,...,\rho}$ according to a discrete uniform distribution.
-This randomized depth-first search traverses the neighbors of each node in ${\mathcal{G}}(v_{0})$ in a random order, which directly corresponds to the sampling process described in line 10 of Algorithm 1.
-Specifically, at each step, we sample the next node in the path from a discrete uniform distribution over the current node’s neighbors within the local subgraph, expressed as $\sim(\mathcal{D}([v^{\prime}\mid(v_{i},v^{\prime})\in{\mathcal{E}}\land v^{\prime}\in{\mathcal{G}}(v_{0})]))$, where $v_{i}$ is the current node in the path.
-To ensure well-defined queries with unique answers, we introduce an additional constraint on path generation.
-This constraint requires that each generated path be unique in terms of its sequence of relationships.
-Specifically, traversing the path from the initial node using the specified relations must lead to a single, unambiguous answer node.
-This approach prevents queries with multiple valid answers, which would complicate the evaluation of the language model’s performance.
+D.3 PrimeKG
+PrimeKG presents a more streamlined structure:
+  • Nodes: The graph consists of nodes representing entities, but unlike Wikidata5m, nodes do not contain accompanying descriptive text.
+  • Edges: Similar to Wikidata5m, edges denote relationships between entities.
+  • Aliases: Each node and edge is associated with exactly one alias, providing a single alternative name rather than multiple variants.
+D.3.1 Preprocessing the PrimeKG knowledge graph
+The PrimeKG dataset requires several key transformations to align with our experimental framework.
+  1. 1.
+    Converting to a Directed Graph: We convert PrimeKG’s undirected structure to a directed graph by creating bidirectional edges.
+Each original edge generates two directed edges with labels incorporating the target entity’s type (e.g., an undirected indication edge between a drug and disease yields ”indication (drug)” and ”indication (disease)” directed edges).
+  2. 2.
+    Enriching Entity Representations:We augment entity representations by prepending type information to each entity’s alias, maintaining a single alias per node to preserve medical precision.
+  3. 3.
+    Generating Supporting Text: To address PrimeKG’s lack of contextual information, we generate descriptive text for each edge using templates tailored to the entity types and relationships (e.g., ”entity_1 is relation name for entity_2”).
+Implementation details are available in the accompanying code.
+  4. 4.
+    Standardizing Character Encoding: All text is converted to ASCII format, matching the Wikidata5m preprocessing pipeline.
+D.4 Query Generation
+Query generation follows a two-phase process: (1) DAG generation according to the knowledge graph structure and specifications, and (2) transformation of the DAG into a natural language query.
+The specifics differ between Wikidata5m and PrimeKG due to their distinct structures and certification requirements.
+D.4.1 Wikidata5m
+For Wikidata5m, we generate linear DAGs (paths) from specified pivot nodes.
+This approach enables focused testing of an LLM’s knowledge comprehension capabilities through a chain of relationships.
+In this case therefore ${\mathcal{N}}_{source}$ has only one node the pivot node, $v_{0}$.
+Then we simulate the algorithm 2 and sample a path for our query generation.
+From the pivot node $v_{0}$, we construct a local subgraph ${\mathcal{G}}(v_{0})$ containing all paths ${\mathcal{N}}_{v_{0}}$ originating from $v_{0}$.
+This local subgraph construction serves two purposes: efficient path sampling and enforcement of the $\rho$ constraint from Algorithm 1.
+We set $\rho=5$ as longer paths often lose semantic coherence in knowledge comprehension queries.
+Within ${\mathcal{G}}(v_{0})$, we generate paths through a randomized sampling process:
+1. Sample path length $k_{choice}$ uniformly from $\{1,\ldots,\rho\}$, following line 1 in Algorithm 2
+2. Generate path through iterative node sampling: - Start with ${\mathcal{N}}=[v_{0}]$ At each step $i$, sample next node according to distribution $\mathcal{D}$ (defined in line 1):
+$v_{i+1}\sim\mathcal{D}([v^{\prime}\mid(v_{i},v^{\prime})\in{\mathcal{E}}\land v^{\prime}\in{\mathcal{G}}(v_{0})])$
+where $v_{i}$ is the current terminal node of ${\mathcal{N}}$
+To ensure well-defined queries per Definition 2.1, we require path uniqueness: following the relation sequence of ${\mathcal{N}}$ from any node must lead to at most one target node.
+This constraint ensures $|{\mathcal{N}}^{\prime}_{sink}|=1$ as specified in Definition 2.1, while preserving the natural graph structure where nodes may have multiple edges with the same relation to non-path nodes.
+Additionally, this approach prevents queries with multiple valid answers, which would complicate the evaluation of the language model’s performance.
 It’s important to note that this uniqueness constraint applies only to the specific path being generated.
@@
 This allowance maintains the natural complexity of the knowledge graph structure, where entities can have multiple relationships of the same type with different entities.
-The pseudocode for the path generation algorithm is specified in  2
+The complete path generation process is detailed in Algorithm 4, which implements the path sampling component of sampleDAG used in lines 2-3 of Algorithm 2.
+The pseudocode for the path generation algorithm is specified in  4
 Algorithm 4 Random Path Generation
@@
 11:  Return $path$
-B.2.2 Query Formulation
-Once a valid path $\Pi$ is generated, we convert it into a query string.
-This process aligns with line 11 in Algorithm 1.
+Query Formulation:
+Once a valid path ${\mathcal{N}}$ is generated, we convert it into a query string.
+This process aligns with line 3 in Algorithm 1.
 The query is constructed by sampling aliases for each node and relation in the path.
-For example, a path $\Pi=[A,B,C]$ might be converted to a query ”sampled_alias(A) $\rightarrow$ sampled_alias((A, B)) $\rightarrow$ sampled_alias(B) $\rightarrow$ sampled_alias((B, C)) $\rightarrow$ ?”.
+For example, a path ${\mathcal{N}}=[A,B,C]$ might be converted to a query ”sampled_alias(A) $\rightarrow$ sampled_alias((A, B)) $\rightarrow$ sampled_alias(B) $\rightarrow$ sampled_alias((B, C)) $\rightarrow$ ?”.
 Here the tuple of two nodes represents their edge.
 The aliases are sampled randomly from a discrete uniform distribution over the available aliases for a node or an edge.
-B.2.3 Example Query Generation
-To illustrate our query generation process, consider the scenario of a path in our subgraph as shown in  9.
+Example Query Generation
+To illustrate our query generation process, consider the scenario of a path in our subgraph as shown in  6.
 Figure 6: Potential Path in a Subgraph where Pivot is Chandler Bing
-Our algorithm would construct the following query from the path presented in  9:
+Our algorithm would construct the following query from the path presented in  6:
 […]
@@
 This example demonstrates how our query generation process creates questions that require multi-hop reasoning, leveraging the structure and relationships within the knowledge graph.
-B.3 Prompt Construction
+D.4.2 PrimeKG
+For PrimeKG, each certificate is defined by a reference DAG ${\mathcal{G}}_{ref}=({\mathcal{N}}_{ref},{\mathcal{E}}_{ref})$ that encodes a specific clinical reasoning pattern.
+Following Algorithm 3, we generate isomorphic DAGs ${\mathcal{G}}^{\prime}$ that preserve the structure and relation types of ${\mathcal{G}}_{ref}$ while varying the specific entities.
+Formally, for a given ${\mathcal{G}}_{ref}$, we sample DAGs ${\mathcal{G}}^{\prime}$ such that there exists a bijection $\phi:{\mathcal{N}}_{ref}\rightarrow{\mathcal{N}}^{\prime}$ preserving: 1.
+Edge structure: $(v_{1},v_{2})\in{\mathcal{E}}_{ref}\iff(\phi(v_{1}),\phi(v_{2}))\in{\mathcal{E}}^{\prime}$ 2.
+Relation types: $(v_{1},v_{2})_{\mathcal{A}}=(\phi(v_{1}),\phi(v_{2}))_{\mathcal{A}}$
+We define nine core reference patterns capturing common clinical reasoning chains:
+  • Which drug is a contraindication for #disease A# and indication for #disease B#, when patient has both diseases?
+  • Which drug can be used as indication for ¡disease A¿ and is atleast an off-label use drug for #disease B#?
+  • Which drug can be used as an indication for #disease A# and is an indication for #disease B#?
+  • Which drug is an indication for #disease A# and interacts synergistically with the treatment of #disease B#?
+  • Which drug targets #genes/proteins(RIDR33)# associated with #disease(RIDR38)#?
+  • Which disease is contraindicated by drug indication #drug A# for #disease A#, which both show #effect/phenotype A#?
+  • Which disease is treated with #drug A# but contraindicated with #drug B#?
+  • Which drug can be used to treat a disease caused by exposure of #exposure A#?
+  • Which disease can be a diagnosis for the following phenotypes - #list of phenotypes#?
+For each reference DAG ${\mathcal{G}}_{ref}$, we implement specialized sampling functions that generate isomorphic DAGs while respecting PrimeKG’s medical domain constraints.
+These functions implement the generateIsomorphisms procedure from Algorithm 3 as follows:
+1.
+Type constraints: Each node $v\in{\mathcal{N}}_{ref}$ specifies required entity types (e.g., DRUG, DISEASE) 2.
+Relation constraints: Each edge $e\in{\mathcal{E}}_{ref}$ requires specific medical relationships (e.g., indication, contraindication) 3.
+Clinical validity: Additional constraints ensure sampled DAGs represent valid clinical scenarios
+For implementation details of these specialized sampling functions, see the discovery_functions in our codebase.
+Query Formulation:
+For a sampled DAG ${\mathcal{G}}^{\prime}$, we generate natural language queries using predefined templates $\tau$ (line 4 in Algorithm 1).
+Each certificate has a minimum of 5 semantically equivalent templates that vary in:
+1.
+Syntactic structure: ”Which drug treats X and Y?” vs.
+”Find a medication indicated for both X and Y” 2.
+Clinical terminology: ”contraindication” vs.
+”should not be used with” 3.
+Question framing: Direct vs.
+scenario-based queries
+D.5 Prompt Construction
 The final prompt is constructed using a template applied to the query. This process involves several steps, each addressing specific requirements:
@@
 These inclusions ensure that the prompt is comprehensive, fits within model constraints, and provides sufficient guidance for the LLM to generate accurate responses.
-B.3.1 Distractor Selection
-Distractors are crucial in assessing an LLM’s comprehension and reasoning abilities.
-We hypothesize that distractors to nodes later in the path closer to the answer would provide more difficulty for the LLM due to their proximity to the answer node.
-Our distractor selection process, implemented in Algorithm 4, employs a weighted sampling approach to prioritize distractors associated with entities closer to the path’s end.
-The algorithm first identifies all potential distractors for each node in the path, then assigns weights inversely proportional to their distance from the final answer node.
-This weighting scheme favors distractors linked to nodes near the path’s end, but still allows for the selection of distractors related to any node in the path.
-This refers to the sampling procedure in the line 12 of the probabilistic progam specification  1.
-By performing weighted sampling from this pool, we ensure a balance between highly relevant distractors and a diverse selection across the entire path.
-B.3.2 Answer Options
+We also provide information on the aliases used and the entities they correspond to in the prompt, to ensure that the LLM knows about the alias.
+D.5.1 Distractor Selection
+Following Definition 2.2, we implement distinct distractor selection strategies for each knowledge graph structure to create challenging evaluation scenarios.
+Wikidata5m Linear Paths
+For a path ${\mathcal{N}}=[v_{1},...,v_{k}]$, we select distractors through weighted sampling based on path position:
+1.
+Distractor Identification: For each node $v_{i}\in{\mathcal{N}}$: - Find candidate distractors $D_{i}=\{\tilde{v}\mid(v_{i},\tilde{v})\in{\mathcal{E}}\land\tilde{v}\notin{\mathcal{N}}\}$ Filter to preserve relation type: $D_{i}^{\prime}=\{\tilde{v}\in D_{i}\mid\exists j>i:(v_{i},\tilde{v})_{\mathcal{A}}=(v_{i},v_{j})_{\mathcal{A}}\}$
+2. Position-based Weighting: Assign weights to $\tilde{v}\in D_{i}^{\prime}$:
+$w(\tilde{v})=i/|{\mathcal{N}}|$
+This weights distractors near path end higher, as formalized in Algorithm 6.
+PrimeKG DAG Structure
+For reference DAG ${\mathcal{G}}_{ref}$ and its isomorphic instance ${\mathcal{G}}^{\prime}$, we:
+1. Include all valid distractors in context:
+[…]
+2. Ensure answer options contain at least one distractor:
+$\text{options}=\{{\mathcal{N}}^{\prime}_{sink}\}\cup\{d\sim\text{Uniform}(D)\}\cup\text{Others}$
+This flexible approach demonstrates how our framework is easily adaptable to different kinds of queries.
+D.5.2 Answer Options
 After formulating the query, we generate a set of answer options. This set includes:
@@
 The process of generating answer options is detailed in Algorithm 5.
-In the algorithm, we sample answer options from the set described above so we are basically sampling from the nodes as in the probabilistic program specifiction line 12  1.
+In the algorithm, we sample answer options from the set described above.
 The answer option algorithm assumes that distractors are input in a list according to the order of preference.
+Another important note is that for primeKG, we try to ensure that all options are of the same type as the correct answer to prevent the LLM from easily eliminating wrong options.
 Algorithm 5 Generate Answer Options
@@
 Algorithm 6 Get Best Distractor
-1:  Input: Graph $G$, Path $\Pi$
+1:  Input: Graph $G$, Path ${\mathcal{N}}$
 2:  Output: $best\_distractor$
@@
 4:  $W\leftarrow$ [] {Weights for distractors}
-5:  for $i\leftarrow$ 00 to len($\Pi$) $-2$ do
-6:     $v\leftarrow\Pi[i]$
+5:  for $i\leftarrow$ 00 to len(${\mathcal{N}}$) $-2$ do
+6:     $v\leftarrow{\mathcal{N}}[i]$
 7:     $N\leftarrow$ GetNeighbors($G$, $v$)
-8:     $N\_distractors\leftarrow$ FilterDistractors($N$, $v$, $\Pi$)
+8:     $N\_distractors\leftarrow$ FilterDistractors($N$, $v$, ${\mathcal{N}}$)
 9:     Extend $D$ with $N\_distractors$
@@
 16:  end if
-B.4 Context Trimming
+D.6 Context Trimming
 To address the input length limitations of various LLMs, we implement a context trimming procedure.
@@
 This involves identifying relevant sentences per edge in the graph and then trimming the context for each query based on this information.
-B.4.1 Finding Relevant Sentences per Edge
+D.6.1 Finding Relevant Sentences per Edge (Wikidata5m)
 Each node in the Wikidata5m knowledge graph has associated textual support for its relations.
@@
     Prepend First Sentence: To ensure the entity’s primary name or common alias is included, we prepend the first sentence of each node’s text to its list of relevant sentences.
-B.4.2 Trimming to Fit Context Length
+D.6.2 Relevant sentences per Edge PrimeKG
+As we synthetically generate data for each edge in PrimeKG, we already have a sentence that corresponds to an edge in primeKG. So no further processing is needed.
+D.6.3 Trimming to Fit Context Length
 When constructing the final prompt for the LLM, we prioritize including the most relevant information within the model’s context length limit.
@@
 12:  return $C$ as $C_{trimmed}$
+Vanilla Setting Note:
+In the vanilla setting, we only want to check simple knowledge comprehension abilities of LLMs, so if our DAG used to create a query has an edge type (alias), we ensure that the provided context trims any sentences which are in the relevant sentences of the same edge type has any edge type in the DAG.
+In distractor setting, there is no such constraint.
 […]
 40
-B.5 Few-Shot Examples
-To guide the LLM towards the desired response format and demonstrate the reasoning process, we include 2 few-shot examples in the prompt.
-These examples provide a clear illustration of how to approach the multi-hop reasoning task.
-We use the following few-shot examples:
+D.7 Few-Shot Examples
+To guide the LLM towards the desired response format and demonstrate the reasoning process, we include 2 few-shot examples in the prompt (common across all prompts to all models).
+These examples provide a clear illustration of how to approach the knowledge comprehension task.
+We investigate the impact of varying number of few-shot examples on LLM performance in Appendix E.
+We use the following few-shot examples for wikidata5m:
 Few Shot Examples Common Context: entity_B is the son of entity_A.
@@
 entity_E Explanation: entity_B $\rightarrow$ (chief of) entity_C $\rightarrow$ (constitutes) entity_D $\rightarrow$ (companion of) entity_E How to get answer: Find what entity_B is the chief of to get entity_C, find what entity_C constitutes to get entity_D, then find the companion of entity_D to get entity_E.
-B.6 Final Prompt
-The final prompt presented to the LLM is constructed using a template that incorporates several key elements:
-Trimmed Context [B.4]: The relevant context extracted and trimmed.
-Query [B.2]: The multi-hop query.
-Answer Options [B.3.2]: The generated answer options, including the correct answer and distractors.
-Few-Shot Examples [B.5]: A set of examples demonstrating the desired response format.
+We use the following few-shot examples for primeKG:
+Few Shot Examples Common Context: (drug) entity_A has side effects: effect_1, effect_5.
+(disease) entity_C has contraindication relation to (drug) entity_E.
+(disease) entity_C is indicated for (drug) entity_A.
+(drug) entity_D has side effects: effect_2, effect_3, effect_4.
+(drug) entity_D has indication for (disease) entity_C.
+(drug) entity_E has indication to (disease) entity_B.
+(gene/protein) entity_F is carrier for (drug) entity_E.
+(disease) entity_B is indicated for (drug) entity_G.
+Question 1: Which drug $\rightarrow$ (has indication for) (disease) entity_C and has $\rightarrow$ (side effect) effect_5 ?
+Options: 1.
+entity_F, n 2.
+entity_E, n 3.
+entity_D, n 4.
+entity_A, n 5.
+entity_B Answer: 2.
+entity_A Explanation: entity_A has the side effects effect_5 while it is the indication of entity_C, so answer is entity_A.
+How to get answer: Find the drugs entity_C is indicated for, which gives us entity_A and then check which of them has the side effect_5.
+Question 2: What drugs should be avoided in (disease) entity_C but are indicated for (disease) entity_B?
+Options: 1.
+entity_F, n 2.
+entity_A, n 3.
+entity_D, n 4.
+entity_E, n 5.
+entity_B Answer: 4.
+entity_E Explanation: entity_B is indicated for entity_E (drug), and entity_E has contraindication to entity_C.
+How to get answer: find the drugs entity_B is indicated for to get entity_E, and entity_G.
+then find the contraindications of entity_C to get entity_E.
+The intersection of these two sets gives the answer, which in this case is entity_E.
+D.8 Final Prompt
+The final prompt presented to the LLM is constructed using a template (same for wikidata5m and primeKG) that incorporates several key elements:
+Trimmed Context [D.6]: The relevant context extracted and trimmed.
+Query [D.4]: The multi-hop query.
+Answer Options [D.5.2]: The generated answer options, including the correct answer and distractors.
+Few-Shot Examples [D.7]: A set of examples demonstrating the desired response format.
 The prompt template is structured as follows:
@@
 The actual number of unique prompts is likely significantly larger, making exhaustive enumeration of all possible generated prompts infeasible.
-B.7 Response Checker function
+D.9 Response Checker function
 We implement a simple response checker function to evaluate the correctness of the model’s answers.
@@
 11:  return $is\_correct$
+Appendix E Ablations
+E.1 Few Shot Prompts
+We conduct an ablation study to examine the impact of varying the number of few-shot examples on Gemini-Flash’s performance in the vanilla task setting.
+While our default configuration uses two few-shot examples, we extend this analysis to include up to five examples.
+Interestingly, we observe no significant variation in performance across these different few-shot configurations.
+The results are presented below in  2.
+Table 2: Certification results for LLMs in vanilla setting with different number of few-shot examples
+Model Avg. lower bound Avg. upper bound Avg. accuracy
+Gemini-1.5-Flash 2Shot (Default) $0.46\pm 0.06$ $0.58\pm 0.06$ $0.52\pm 0.06$
+Gemini-1.5-Flash 3Shot $0.46\pm 0.06$ $0.58\pm 0.06$ $0.52\pm 0.06$
+Gemini-1.5-Flash 4Shot $0.46\pm 0.07$ $0.58\pm 0.07$ $0.52\pm 0.07$
+Gemini-1.5-Flash 5Shot $0.46\pm 0.07$ $0.58\pm 0.07$ $0.52\pm 0.07$
+E.2 Distractor Distributions
+To assess the impact of distractor distribution on model performance, we implement three distinct distractor distribution strategies:
+  1. 1.
+    Tail-weighted: Linearly increasing weights towards the tail end of the path, prioritizing distractors near the answer node. This serves as our default setting.
+  2. 2.
+    Head-weighted: Linearly increasing weights towards the head of the path, emphasizing distractors near the query’s starting point.
+  3. 3.
+    Uniform: Equal probability of selecting distractors from any position along the path.
+We observe no significant differences in either of the settings. The results are presented in  3 below.
+Table 3: Certification results for Gemini-Flash with different distractor distributions
+Model Avg. lower bound Avg. upper bound Avg. accuracy
+Gemini-1.5-Flash Setting 1 (Default) $0.42\pm 0.10$ $0.55\pm 0.10$ $0.48\pm 0.10$
+Gemini-1.5-Flash Setting 2 $0.42\pm 0.11$ $0.55\pm 0.11$ $0.48\pm 0.11$
+Gemini-1.5-Flash Setting 3 $0.42\pm 0.11$ $0.55\pm 0.11$ $0.48\pm 0.11$
+E.3 Model Performances with varying Path Length
+Among our certificates, we have queries of various lengths.
+Here we study the effects on models behavior on queries with varying length by considering the number of hops they require to reason to answer the query(which is 1 less than the path length).
+To do so, we refer to the number of hops to answer a query as k where $1\leq$ k $<\rho$.
+Varying Setting: In figure  7 we show plots for various specifications for the GPT4o model.
+Figure 7: Variations in the bounds against the path lengths across various specifications.
+Varying Quantization: In figure  8 we show plots when the quantization is varied with the Llama3-8B model on the shuffle specification and their effects on performance.
+Figure 8: Variations in the bounds against the path lengths across various quantizations.
+Varying Models: In fig  9 we show plots for the shuffle specification and performance across the models(the open-source models use fp16 precision).
+Figure 9: Variations in the bounds against the path lengths across various models in the shuffle setting.
+In Figure  7, we observe that the performance across settings converges as k increases and the distractor setting is most impactful on the performance for $k=2$.
+In Figure  8, we infer that as k increases the performance of the models’ on the task converges across the different quantizations.
+We hypothesize this is due to the increasing complexity of the reasoning task.
+In Figure 9, we see that larger models (GPT-4o, Gemini-Pro) show less severe drop in performance compared to their smaller models.
+The figure shows that large models may have learnt to better apply 1-step reasoning for multiple steps when compared to their smaller counterparts.
+E.4 Chain of Thought Prompting
+We also conduct an ablation on how Chain-of-Thought(COT) prompting can affect the performance of language models on the knowledge comprehension task.
+Specifically, we investigate the Phi-3 3B model (precision: float16) in the vanilla setting with COT prompting strategy.
+We augmented our standard few-shot examples (D.7) with COT steps and added structured reasoning guidance to the prompt template (D.5):
+COT additions to prompt template Answer in the following the below format: Let’s solve this step by step: 1) Let’s identify the starting point and path: - Start: [identify starting entity] - Path to follow: [break down the path components] 2) Let’s follow the path: Starting from [entity] → [first relationship] → [next entity] → [next relationship] → [next entity] … [continue as needed] 3) Verify our final destination reaches one of the given options Therefore, the correct answer is: $<$option_number$>$.
+$<$option_text$>$
+In the vanilla setting, adding COT prompting improved Phi-3 3B’s performance, with the bounds increasing by 0.11 summarized in Table 4.
+While we acknowledge the potential benefits of COT, earlier experiments were limited due to the significantly increased computational cost (generating 5-8 times more tokens) and the expenses of COT, particularly with closed-source models as output tokens are much more expensive.
+Table 4: Certification results for Phi-3 3B with and without COT
+Prompting Strategy Avg. lower bound Avg. upper bound Avg. accuracy
+No COT (Default) $0.34\pm 0.05$ $0.46\pm 0.06$ $0.40\pm 0.05$
+COT $0.45\pm 0.08$ $0.57\pm 0.08$ $0.51\pm 0.08$
+Appendix F Case Studies from certifying over Wikidata5m
+We analyze the certification results, qualitatively.
+First, we show the responses of $3$ models in Figure 10 — Phi-3 (3B), GPT-4o, and Gemini, obtained when certifying them for the Vanilla specification defined over a subgraph pivoted at the node for ‘Batman Begins’ movie.
+The samples reflect the certificates.
+Next, we identify and categorize prominent kinds of model responses.
+We frequently see the following failure modes — distracted and missed relation.
+In the former, the model gets deviated from the query by following the distractor context in its prompt, resulting in an incorrect answer.
+In the latter, the model skips some reasoning steps needed for the final correct answer.
+In cases of good reasoning, model accurately follows the query and gives the correct answer.
+Figure 11 presents examples of the aforementioned kinds of model responses for GPT-4o.
+Figure 10: Qualitative analysis of samples used for certifying knowledge comprehension for Vanilla specifications on the Wikidata5m subgraph pivoted at the node for ‘Batman Begins’ movie.
+The context provided in the prompts is not shown for brevity.
+Wrong model responses are colored red and correct ones are colored green.
+The samples are consistent with our results, wherein Phi-3 (3B) has lower certification bounds than GPT-4o’s bounds, which are lower than those for Gemini-Pro.
+Figure 11: GPT-4o responses showing good reasoning, failures due to distractors or skipped reasoning